### v2

# ScholaWrite: A Dataset of End-to-End Scholarly Writing Process

Abstract
Writing is a cognitively demanding task involving continuous decision-making, heavy use of working memory, and frequent switching between multiple activities. Scholarly writing is particularly complex as it requires authors to coordinate many pieces of multiform knowledge. To fully understand writers’ cognitive thought process, one should fully decode the end-to-end writing data (from individual ideas to final manuscript) and understand their complex cognitive mechanisms in scholarly writing. We introduce ScholaWrite dataset, the first-of-its-kind keystroke logs of an end-to-end scholarly writing process for complete manuscripts, with thorough annotations of cognitive writing intentions behind each keystroke. Our dataset includes LaTeX-based keystroke data from five preprints with nearly 62K total text changes and annotations across 4 months of paper writing. ScholaWrite shows promising usability and applications (e.g., iterative self-writing) for the future development of AI writing assistants for academic research, which necessitate complex methods beyond LLM prompting. Our experiments clearly demonstrated the importance of collection of end-to-end writing data, rather than the final manuscript, for the development of future writing assistants to support the cognitive thinking process of scientists. Our de-identified dataset, demo, and code repository are available on our project page111https://minnesotanlp.github.io/scholawrite/.
ScholaWrite: A Dataset of End-to-End Scholarly Writing Process
Linghe Wang††thanks: Equal contribution. Minhwa Lee Ross Volkov Luan Tuyen Chau Dongyeop Kang University of Minnesota {wang9257,lee03533,volko032,chau0139,dongyeop}@umn.edu
1 Introduction
Scholarly writing requires researchers to produce texts that concisely and precisely present novel findings and to follow a systematized structure and style by their targeted venue Jourdan et al. (2023); Kallestinova (2011); Bourekkache (2022). To address this need, researchers have leveraged large language models (LLMs) to develop intelligent support systems in several writing tasks, such as revision process Du et al. (2022b); Kim et al. (2022) or feedback generation Liang et al. (2024); Lu et al. (2024a). However, a critical question arises: Are LLMs able to truly understand and generate writing as human scientists do?
LLMs are generally trained to progress autoregressively, generating text from left to right. In contrast, human writing typically involves multiple iterations of complex and non-linear cognitive actions to refine the main message Flower and Hayes (1981); Koo et al. (2023), as illustrated in Figure 1. This fundamental cognitive gap between human writing and language models necessitates a deeper understanding of human writing processes and the development of cognitively-aligned writing assistance tools, moving beyond basic auto-completion.
Although much literature has examined the cognitive processes used in general writing tasks Lindgren and Sullivan (2003); Hayes and Berninger (2014); Zhu et al. (2023); Wengelin and Johansson (2023), the specific cognitive patterns in scholarly writing remain understudied. Our work observes the end-to-end writing process of scholarly writing, inspired by the method of keystroke collections that have been a major methodology to observe individual writing processes Baaijen et al. (2012); Latif (2008); Leijten and Van Waes (2013) in cognitive science or a few recent works in NLP at a small scale Koo et al. (2023); Velentzas et al. (2024).
We believe our work (ScholaWrite) is the first attempt to present a keystroke corpus of scholarly writing with annotations of cognitive writing processes, which were collected over multiple months and produced by early-career researchers. Furthermore, we developed novel data collection and annotation systems using Chrome extension, which improves data accessibility to the public and enables real-time collection of LaTeX-based keystrokes from multiple researchers. We also present a comprehensive taxonomy of cognitive writing processes specific to the scholarly writing domain. Throughout several experiments with LLMs, ScholaWrite shows promising usability as a resource for enabling LLMs to understand the human cognitive process of scholarly writing and provide writing suggestions that are aware of writers’ cognitive behaviors.
Contributions
We present ScholaWrite, a curated dataset of nearly 62K LaTeX-based keystrokes that were turned into publications in the computer science domain, annotated by experts in linguistics and computer science. We develop a taxonomy of scholarly writing intentions, providing an overall understanding of how scholars tend to develop writing manuscripts. Experiment results showed that the Llama-8B model fine-tuned on the ScholaWrite dataset achieved the high linguistic quality of the final writing.
2 Related Work
2.1 Cognitive Theory of Writing Process
Research on human writing significantly shifted from focusing solely on the final written product to exploring the cognitive stages of various writing phases Diederich (1974); Perl (1979); Krapels (1990); MacArthur and Graham (2016).
Building on this process-oriented approach, Flower and Hayes (1981)’s cognitive writing theory outlines three key cognitive sub-processes grounded in empirical evidence: (1) planning; (2) translating; and (3) reviewing. These processes are complex, interconnected, and non-linear, with writers moving fluidly between them. Our work expands this model to develop our own taxonomy of more granular cognitive patterns of human writing, specifically in scholarly communication.
Koo et al. (2023) presented a taxonomy of the scholarly writing process based on keystroke collections of just a short research plan over a thirty-minute session. Inspired by this work, we extend their approach to a larger-scale collection of keystrokes over months, whose final written products turned into research publications. Reviewed by a linguistics expert, our comprehensive taxonomy is built on top of prior literature and encompasses end-to-end writing trajectories of the cognitive process in a scholarly domain.
2.2 Keystroke Loggers for Scholarly Writing
Keystroke logging tools such as Inputlog allow researchers to observe digital writing without interfering with the writer’s process Chan (2017); Johansson et al. (2010); Leijten and Van Waes (2013); Lindgren and Sullivan (2019). However, they tend to operate in a closed ecosystem, i.e., only running in MS Word. This makes the current tools less accessible, especially for the scientific communities who use other typesetting such as LaTeX. Moreover, it is not well-suited for collecting data from extended writing sessions.
Currently, such keystroke logging studies have rarely examined scholarly communication settings that use LaTeX, especially for scientific writing purposes. To address this gap, our work implemented novel several interconnected systems that securely collect keystrokes in real-time over months from Overleaf, a widely-used online LaTeX editor, store them without privacy issues, and retrieve them for annotating writing intentions at span level. Our system workflow provides a more natural environment for observing cognitive writing activities by multiple researchers over long periods and across multiple sections without interrupt their writing.
2.3 Datasets for Scholarly Writing
Previous work primarily focused on constructing datasets for understanding the writing process of scholarly domains, which vary in their content and scale. Much of the publicly available datasets analyzed linguistic style changes or grammatical edits during the revision process Du et al. (2022b); Jiang et al. (2022); Ito et al. (2019); Mita et al. (2022). Also, the datasets from D’Arcy et al. (2024); Jourdan et al. (2024); Kuznetsov et al. (2022) collected edits conditioned on feedback and peer review process, while Kobayashi et al. (2022); Narimatsu et al. (2021) constructed datasets for citation generations. Existing work is also limited to only specific sections of scientific papers (e.g., abstract, introduction) Du et al. (2022b); Mita et al. (2022). In comparison, we construct a dataset that covers all cognitive phases of the writing process from its beginning to the final written product, through long periods of the data collection process.
Furthermore, those existing corpora for scientific writing research are created from the comparison between multiple versions of full-length, final products retrieved from preprint databases Jiang et al. (2022); Du et al. (2022b); D’Arcy et al. (2024); Jourdan et al. (2024). As pointed out by Jourdan et al. (2023), the comparison of finalized outputs does not present how those manuscripts evolved. Our work addresses this problem by building a corpus of keystrokes each of which represents the progress of the publications (See Figure 2).
3 Dataset Construction
To preserve writers’ natural thought flow and collect authentic writing data, we developed a Chrome extension that records keystroke data in the background. After the study, an in-depth annotation process is conducted to interpret the intent behind each keystroke. For open research, we make all the code of our data collection and annotation framework publicly available: https://minnesotanlp.github.io/scholawrite/. We describe the detailed process of our data collection, annotation system design and taxonomy creation in the following sections.
3.1 Participant Recruitment
We recruited 10 graduate students222All of them currently attend a four-year university in the U.S. and are proficient in English. from the computer science department who were actively preparing academic manuscripts in Overleaf LaTeX editor333https://overleaf.com/ for submission to peer-reviewed conferences in AI and NLP. The data is collected from November 2023 to February 2024, totaling up to 4 months. Our study has been approved by the IRB institution of the authors. Please refer to Appendix A.1 for more detailed descriptions.
3.2 Chrome Extension for Keystroke Collection in Overleaf
We designed and implemented a Chrome extension (Figure 4(a)(a)), which enables the real-time collection of keystroke trajectories in the Overleaf platform. Participants can create their account credentials, and after logging into the system, the extension monitors the user’s keystrokes in the background silently, without disrupting the typical writing process.
When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel. When each of these actions444Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page. occurs, the extension uses ‘diff_match_patch’ package555https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 3). Then, the extension will send the difference array along with other metadata (e.g., time stamp, author ID, file name, etc.) to the backend server. To prevent any issue of private data collection, the backend fetches only the IDs of the Overleaf projects that participants consented to share during the recruitment process and filters out participants’ keystroke data from any unauthorized projects666We used Google Sheet API to retrieve ID information that we collected during the recruitment process..
For any Overleaf project that consists of multiple LaTex files, which is common in collaborative scholarly writing, we also collected all keystrokes from subfiles associated with the main LaTeX file. Our comprehensive data collection process captures the end-to-end writing processes of the participating authors across all parts of Overleaf projects. This approach ensures that our dataset reflects the full scope of scholarly writing including edits made in auxiliary files such as files of each section, appendix, bibliography, etc.
All keystroke data stored in the database were processed for annotation and model training purposes. Here, raw keystrokes from the Chrome extension system are stored as arrays that represent the differences between subsequent keystrokes, along with accompanying metadata.
| 1st | Intention | Definition | An example action | Prop. |
| Planning | Idea Generation | Formulate and record initial thoughts and concepts. | writing down keywords of a paragraph beforehand (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
| Idea Organization | Select the most useful materials and demarcate those generated ideas in a visually formatted way. | Linking the generated ideas into a logical sequence and spacing out between ideas (e.g., “..% (1) need diff. stress testing…%%[Spacing] (2) exp. setup? ”) | 0.5% | |
[…]
| Implementation | Text Production | Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source. | Generating subsequent sentences with the author’s own idea (e.g., “… GPT-4 (OpenAI, 2023) explains the data … Our approach is built on top of GPT-4…”) | 57.4% |
[…]
| Coherence | Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole. | “Each comment was annotated by three different annotators , which and we achieved high inter-annotator agreement.” | 3.3% | |
| Clarity | Improve the semantic relationships between texts to be more straightforward and concise. | “..relevant studies have examined one of the several textual styles one aspect of texts, the formality, ….” | 11.5% | |
| Structural | Improve the flow of information by modifying the location of texts and objects. | “We calculate Pearson’s correlation for human alignment to compare the alignment between lexicon-based preferences and humans’ original preferences. First, we calculated the score from each human participant to compare the alignment between lexicon-based preferences and humans’ original preferences.” | 3.7% | |
| Linguistic Style | Modify texts with the writer’s writing preferences regarding styles and word choices, etc. | “We believe posit that …” | 1.6% | |
| Scientific Accuracy | Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims. | “..Pearson’s correlation (0.780.68; p < 0.01)” | 0.7% | |
[…]
3.3 Writing Intention Annotation
To capture comprehensive writing behaviors, we designed a data collection process that involves running a Chrome extension in the background. However, keystroke data alone may be insufficient to fully understand the nuanced intentions behind each writing action. To bridge this gap, we implement an additional annotation step to decode the writer’s intent for each keystroke using a specialized annotation interface. This process is crucial for developing a detailed taxonomy of the end-to-end scholarly writing process, which serves as the basis for annotating the collected keystrokes. Further details about the annotation process can be found in Appendix B.
To perform annotations, we first processed those raw keystroke entries by file name, type of writing actions (e.g., add, delete), an array of differences between two subsequent texts, and line number in an Overleaf editor.
Annotation Interface
We developed a novel interface for the keystroke annotation process. The interface offers various modes for visualizing keystroke collections within each Overleaf project: by time, LaTeX file, and author (Figure 4(a)(b)).
In the right panel, annotators can interactively right-click to navigate the timeline of collected keystrokes in different colors (e.g., red for deletions and green for additions). They then identify a range of keystrokes whose underlying purpose or goal is similar when making edits during the writing process (e.g., generating outlines or correcting grammatical errors). They can manually enter the beginning and ending indices and annotate the span by selecting a pre-defined label of the relevant intention from the dropdown menu in the left panel. Furthermore, the interface shows the metadata information of each keystroke, along with the selected label.
Intention Taxonomy Construction
Our taxonomy (Table 1) builds on prior literature Flower and Hayes (1981); Du et al. (2022b); Koo et al. (2023). During annotations, we have identified 15 types of scholarly writing intentions from the five projects and further grouped them into three high-level categories according to Flower and Hayes (1981) - Planning, Implementation, and Revision.
Following Pustejovsky et al. (2017)’s iterative annotate-and-refine framework, we performed an iterative open coding for the collected keystrokes from the two Overleaf projects. Using the annotation interface, each annotator reviewed the first few writing keystrokes and identified the high-level processes based on (Flower and Hayes, 1981)’s initial taxonomy of cognitive writing (planning, implementation, and revision). Next, the annotators (1) determine the span of keystrokes that form a meaningful unit together (e.g., a clause or sentence) within the high-level process and (2) decide the underlying intention of edits. The annotators coded these spans independently, resolving conflicts through multiple in-depth discussions. Finally, a cognitive linguist participated in all discussions, refining the definition and scope of each writing intention.
To ensure the validity of our taxonomy, we followed the two standard practices Nickerson et al. (2013); Kundisch et al. (2021): (1) All elements should be mutually exclusive and (2) the taxonomy should be collectively comprehensive. To achieve this, we flagged any ambiguous keystroke spans that could be interpreted as representing multiple intentions. Through iterative discussion, annotators refined these spans’ characteristics and improved each category’s decision boundaries. Once the codebook and taxonomy were finalized, we updated intention annotations777The list of those intentions is then updated in the dropdown bar of the annotation interface. for the selected two projects and applied the taxonomy to the remaining projects. The final inter-annotator agreement is using a weighted F1 score in a multi-label, multi-class setting for a randomly selected window of 1K keystrokes. Please refer to Appendix B for more details about the entire annotation procedures.
4 ScholaWrite: A Dataset of Cognitive Process of Scholarly Writing
| Project | 1 | 2 | 3 | 4 | 5 |
| # Authors | 1 (3) | 1 (4) | 1 (3) | 1 (4) | 9 (18) |
| # keystrokes | 14,217 | 5,059 | 6,641 | 8,348 | 27,239 |
| # Words added | 17,387 | 23,835 | 7,779 | 12,448 | 57,511 |
| # Words deleted | 11,739 | 15,158 | 2,308 | 7,621 | 25,853 |
| 1 | 2 | 3 | 4 | 5 | |
| Idea Generation | 515 | 130 | 116 | 309 | 3255 |
| Idea Organization | 0 | 45 | 25 | 9 | 231 |
| Section Planning | 182 | 57 | 111 | 201 | 773 |
| Text Production | 9267 | 2438 | 5109 | 4478 | 14031 |
| Object Insertion | 583 | 383 | 62 | 486 | 1300 |
| Cross-reference | 141 | 112 | 13 | 292 | 458 |
| Citation Integration | 75 | 151 | 69 | 127 | 245 |
| Macro Insertion | 16 | 7 | 51 | 29 | 33 |
| Linguistic Style | 233 | 75 | 42 | 201 | 411 |
| Coherence | 422 | 242 | 126 | 193 | 1021 |
| Clarity | 1249 | 645 | 721 | 1180 | 3301 |
| Scientific Accuracy | 307 | 15 | 2 | 24 | 95 |
| Structural | 359 | 506 | 105 | 257 | 1042 |
| Fluency | 116 | 90 | 46 | 135 | 476 |
| Visual Formatting | 752 | 163 | 43 | 427 | 567 |
Summary
Our post-processed dataset, ScholaWrite, contains end-to-end writing trajectories with intentions annotated, for five Overleaf projects whose final product turned into arXiv preprints. The dataset consists of 61,504 arrays of keystroke changes with fine-grained annotations of writing intentions. We publicly release our post-processed data in Huggingface data cards.888https://huggingface.co/datasets/minnesotanlp/scholawrite Table 3 shows the overall statistics of the dataset.
Writing Intention Distributions
The writing intention distributions is illustrated in Figure 5(a). During the planning stage, idea generation is the most frequent intention. Text production dominates the implementation stage, and clarity becomes the primary focus during the revision stage. Since we defined the boundary between implementation and revision as the point where meaning changes occur, the high frequency of text production suggests that authors often continue revising their texts by introducing new topics into the documents. Table 3 shows the label distribution per project.
Flows Between Writing Intentions
We analyze the flow of writing intentions most likely to follow each other throughout the writing process, as presented in Table 4. Many intentions proceed to text production, where the idea generation has the highest probability of feeding into text production. Also, the reflexive relationship between text production and clarity suggests a constant loop of producing new texts and refining them, ensuring the precision and reasonability of their claims.
| Label | Subsequent label | Probability | |
| Idea Generation | Text Production | 0.52 | |
| Idea Organization | Idea Generation | 0.34 | |
| Section Planning | Text Production | 0.33 | |
| Text Production | Clarity | 0.20 | |
| Object Insertion | Text Production | 0.32 | |
| Citation Integration | Text Production | 0.37 | |
| Cross-reference | Text Production | 0.36 | |
| Macro Insertion | Idea Generation | 0.29 | |
| Fluency | Text Production | 0.30 | |
| Coherence | Text Production | 0.34 | |
| Clarity | Text Production | 0.35 | |
| Structural | Text Production | 0.27 | |
| Linguistic Style | Text Production | 0.29 | |
| Scientific Accuracy | Text Production | 0.34 | |
| Visual Formatting | Text Production | 0.25 |
Time-series Distributions of Writing Intentions
We calculated the average Wasserstein distance between each intention distribution and uniform distribution to assess how evenly intentions occur throughout the writing process. As shown in Figure 5(b), text production had the smallest distance, while macro insertion had the largest. This suggests that text production is spread throughout the writing, whereas macro insertion is concentrated in specific phases. A similar pattern appears in Figure 5(c), where text production is evenly distributed over time, and macro insertion peaks in the early and final stages of the writing process. Figures 6(a) and 6(b) demonstrate a similar pattern that text production of the implementation stage spreads evenly over time. Types of clarity and structural revisions tend to occur after the medium steps of writing. Please refer to Figures 9 to 12 in the Appendix for each of the five Overleaf projects.
5 Applications for Writing Assistance with ScholaWrite
We believe ScholaWrite can serve as a valuable resource for training language models or enhancing the cognitive writing process in future writing assistants for scholarly writing. To evaluate its usability, we performed several experiments where existing LLMs were trained to mimic the complex, non-linear scientific writing processes of human scholars. Specifically, we aimed to showcase the capabilities of LLMs trained on ScholaWrite in two scenarios:
(1) Predicting an author’s next writing intention (Section 5.1)
: The task is crucial for writing assistants to accurately assess the writer’s current status in context and predict the correct writing intention. This enables them to offer cognitively-appropriate writing suggestions that align with the writer’s needs.
(2) Iteratively generating scholarly writing actions from scratch (called Iterative Self-Writing), mirroring the human writing process (Section 5.2)
: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
As illustrated in Figure 7, the first task (represented by the inner box labeled "Prediction") takes the "before" text from a keystroke pair and context prompt, and predicts the writing intention to apply for the subsequent actions. Once the intention is predicted, the second task generates the “after” text based on the predicted intention and given prompt (represented by the outer box labeled “Generation”). This process repeats iteratively until no further changes are made or a maximum number of iterations, such as 100, is reached.
We randomly split our dataset into train (80%) and test (20%) sets. For each intention label in the test set, we randomly select up to 300 keystroke entries, due to budget constraints. Note that the same train and test sets are applied to all models across all experiments. See Appendix D for a full description of the model training process.
5.1 Predicting Next Writing Intention
Setup & Metrics
We use BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and Llama3.1-8B-Instruct Dubey et al. (2024) as baselines, fine-tuning each on the ScholaWrite training set. For comparison, we also run GPT-4o OpenAI (2024) on the test set. The fine-tuning prompt included all possible labels with definitions, task instructions, the “before” text chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”. Differences in prompts were limited to only task instructions (see Appendix D.3 for prompt details). To evaluate model performance, we used a weighted f-1 score999Weighted F-1 was chosen to address skewed label distribution (as shown in Figure 5(a) and Table 1)., comparing predictions to gold intention labels on the test set.
| BERT | RoBERTa | LLama-8B | GPT-4o | |
| Base | 0.04 | 0.02 | 0.12 | 0.08 |
| + SW | 0.64 | 0.64 | 0.13 | - |
Results
Table 5 presents the weighted F1 scores for predicting writing intentions across baselines and fine-tuned models. Regardless of the intricate nature of the task itself101010Each model is asked to predict the next intention by only giving the ‘before’ text. Human annotators, however, label the data by looking through multiple consecutive ‘before and after’ edits to determine the current intention. In addition, the next intention chosen by the author does not necessarily mean that it is the only correct intention., all models finetuned on ScholaWrite show an improved performance compared to their baselines. BERT and RoBERTa achieved the most improvement, while LLama-8B-Instruct showed a modest improvement after fine-tuning. As detailed in Appendix D, more training epochs and encoder-decoder architectures of BERT variants are assumed to be the reason for significant improvement compared to LLMs. This aligns with findings from Grasso and Locci (2024); Yu et al. (2023), which shows that RoBERTa and BERT can often match or even outperform LLMs for the text classification tasks. Those results demonstrate the effectiveness of our ScholaWrite dataset to align language models with writers’ intentions.
5.2 Iterative Self-Writing
Setups
Iterative self-writing involves two subtasks (Figure 7): (1) next intention prediction and (2) "after-text" generation. For intention prediction, we use fine-tuned Llama3.1-8B-Instruct on the ScholaWrite training set (“Llama-8B-SW-pred”), compared to vanilla models (Llama-8B-Instruct, and GPT-4o) from Section 5.1 as baselines. Please refer to Figure 13 in Appendix D for the detailed pipeline.
For “after-text” generation, another vanilla Llama3.1-8B-Instruct is fine-tuned on the ScholaWrite training set (“Llama-8B-SW-gen”), with baselines including vanilla Llama3.1-8B-Instruct and GPT-4o. The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text. The model’s task is to generate “after-text” given the verbalizer and “before” text. To maintain consistency across outputs, prompts were adjusted for each model, with differences mainly in task instructions (see Appendix D.3 for the prompt templates).
Seed documents were derived from four award-winning NLP papers spanning different topics Zeng et al. (2024); Lu et al. (2024b); Du et al. (2022a); Etxaniz et al. (2024). Listing LABEL:table:seed-entry presents the LaTeX-formatted abstract of Du et al. (2022a) as an example seed text for initiating the iterative self-writing experiment.
6 Conclusion and Future Work
We introduce ScholaWrite, the first-of-its-kind dataset that captures the end-to-end scholarly writing process, consisting of nearly 62K LaTeX keystrokes collected via our custom-built Chrome extension. This dataset, sourced from ten graduate students with varying levels of scientific writing expertise, is further enriched with expert annotations based on a novel taxonomy of cognitive writing intentions inspired by Flower and Hayes (1981); Koo et al. (2023). Our study demonstrates the value of ScholaWrite as a resource for advancing the cognitive capabilities of large language models (LLMs) and developing cognitively-aligned writing assistants, enabling them to emulate the complex, non-linear, and intention-driven nature of human writing.
Future work includes expanding the dataset to diverse academic fields, authors, and collaborative projects, thus enabling models to generalize beyond fact recall to emulate human decision-making and reasoning in more realistic academic environments. Additionally, integrating advanced memory architectures and lifelong learning techniques could further enhance LLMs’ ability to adapt dynamically to evolving writing intentions and produce coherent, high-quality scholarly outputs.
Limitations and Ethical Considerations
We acknowledge several limitations in our study. First, the ScholaWrite dataset is currently limited to the computer science domain, as LaTeX is predominantly used in computer science journals and conferences. This domain-specific focus may restrict the dataset’s generalizability to other scientific disciplines. Future work could address this limitation by collecting keystroke data from a broader range of fields with diverse writing conventions and tools, such as the humanities or biological sciences. For example, students in humanities usually write book-length papers and integrate more sources, so it could affect cognitive complexities.
Second, our dataset includes contributions from only 10 participants, resulting in five final preprints on arXiv. This small-to-medium sample size is partly due to privacy concerns, as the dataset captures raw keystrokes that transparently reflect real-time human reasoning. To mitigate these concerns, we removed all personally identifiable information (PII) during post-processing and obtained full IRB approval for the study’s procedures. However, the highly transparent nature of keystroke data may still have discouraged broader participation. Future studies could explore more robust data collection protocols, such as advanced anonymization or de-identification techniques, to better address privacy concerns and enable larger-scale participation. We also call for community-wise collaboration and participation for our next version of our dataset, ScholaWrite 2.0 and encourage researchers to contact authors for future participation.
Furthermore, all participants were early-career researchers (e.g., PhD students) at an R1 university in the United States. Expanding the dataset to include senior researchers, such as post-doctoral fellows and professors, could offer valuable insights into how writing strategies and revision behaviors evolve with research experience and expertise. Despite these limitations, our study captured an end-to-end writing process for 10 unique authors, resulting in a diverse range of writing styles and revision patterns. The dataset contains approximately 62,000 keystrokes, offering fine-grained insights into the human writing process, including detailed editing and drafting actions over time. While the number of articles is limited, the granularity and volume of the data provide a rich resource for understanding writing behaviors. Prior research has shown that detailed keystroke logs, even from small datasets, can effectively model writing processes Leijten and Van Waes (2013); Guo et al. (2018); Vandermeulen et al. (2023). Unlike studies focused on final outputs, our dataset enables a process-oriented analysis, emphasizing the cognitive and behavioral patterns underlying scholarly writing.
Third, collaborative writing is underrepresented in our dataset, as only one Overleaf project involved multiple authors. This limits our ability to analyze co-authorship dynamics and collaborative writing practices, which are common in scientific writing. Future work should prioritize collecting multi-author projects to better capture these dynamics. Additionally, the dataset is exclusive to English-language writing, which restricts its applicability to multilingual or non-English writing contexts. Expanding to multilingual settings could reveal unique cognitive and linguistic insights into writing across languages.
Fourth, due to computational and cost constraints, we evaluated the usability of the ScholaWrite dataset with a limited number of LLMs and hyperparameter configurations. As shown in Table 5, the Llama-8B-Instruct model demonstrated only marginal improvements after fine-tuning on our dataset. This underscores the need for future research to explore advanced techniques, such as fine-grained prompt engineering, to better align LLM outputs with human writing processes. Specifically, optimizing prompts with clearer contextual guidance (e.g., "before-text" and intention label definitions) may significantly enhance model performance.
Finally, the human evaluation process in Section LABEL:sec:appendix:human-eval was determined as exempt from IRB review by the authors’ primary institution, while the data collection using our Chrome extension program was fully approved by the IRB at our institution. Importantly, no LLMs were used during any stage of the study, except for grammatical error correction in this manuscript.
Acknowledgements
This work was supported by the research gift from Grammarly. We thank Anna Martin-Boyle and Ryan Koo for their insights and the data collection tool based on their earlier version of this paper. We thank the participants of our data collection for their valuable time, effort, and contributions, which were essential to the success of this research. We also thank Minnesota NLP group members and anonymous reviewers for providing us with valuable feedback and comments on the paper draft.
Appendix A More About Data Collection Process
A.1 Participant Recruitment & Demographics
We recruited ten graduate students in the computer science department who actively prepared their manuscripts in Overleaf, an online LaTeX editor, and who aimed to submit their manuscripts to peer-reviewed conferences. We held a consent procedure with each participant through a 30-minute virtual meeting remotely. After the consent process, we installed on their computers a Chrome extension program that we designed and implemented only for this study and asked for the ID number of only the Overleaf projects that the participants agreed to share as their manuscripts. We provided all participants with a $100 Amazon gift card per project, which could be divided among the authors involved in the project. Our data collection process is approved by the IRB of the authors’ primary institution.
All ten participants of our study are graduate students who currently study computer science domain at an accredited university in the United States. Out of the ten, two of them identified themselves as native English speakers, and the remaining participants identified themselves as proficient in English in terms of writing. Also, two of the ten participants attained a Master of Science degree in Computer Science with several publication experiences, and the remaining eight of them are currently PhD students with extensive research experiences.
A.2 Technical Details of System Implementations
We explain the technical implementation details of the two systems for the data collection process. For the Chrome extension, we implemented a backend application using Flask and Python and stored all keystroke data in the MongoDB database.
For the annotation interface, we used HTML/CSS and JavaScript for the client side and Flask for the backend. All data for the annotation interface was retrieved from the MongoDB database used in the Chrome extension system.
A.3 Data Post-Processing
For use during the annotation phase, each keystroke entry from the raw collection includes the following fields: (1) a valid file name; (2) a valid writing action that triggered keystroke logging (e.g., copy, paste, typing, etc.); (3) a valid array of differences to enable visualization of writing trajectories; and (4) the line numbers in the Overleaf editor. Data entries annotated with a valid intention label (i.e., labels except ‘artifact’) and having a difference array length of fewer than or equal to are then used for model training.
Regarding the additional postprocessing for public use, we took the following steps to post-process our data with the annotations to promote the usability of our dataset and prevent any privacy issues. For the annotation data, we only include the keystroke changes, anonymized project ID, timeframe information, and anonymized author’s name (e.g., 0, 1, 2, etc.) from the metadata. Then, we extract the before and after texts from the differences array. We also include the annotated intention label for each entry.
[…]
Appendix B More About Writing Intention Annotations & Taxonomy
B.1 Annotator Recruitment
Due to privacy concerns, we did not hire external freelancers with expertise, rather the two corresponding authors of this paper annotated the data are graduate students who possess extensive scholarly writing experiences in natural language processing and data annotation skills. The raw keystroke data collected by our Chrome extension could potentially contain personally identifiable information, such as specific content edits or metadata that could reveal the identity of the authors. To ensure the confidentiality and ethical handling of sensitive information, we restricted access to the data to the authors only. This annotation process was also authorized by the IRB of the authors’ institution. Please note that the final dataset which will be released publicly is ensured not to contain any PII information through several sophisticated post-processing steps.
B.2 Detailed Annotation Process
The two authors (or annotators) collaborated with a cognitive linguist to develop a codebook and review the results of the annotations. Also, the two annotators conducted an iterative open coding approach to identify several unique writing intentions from keystrokes and developed a codebook of intention labels (“ground-truth labels”) within each high-level process (Planning, Implementation, and Revision) based on the findings from Flower and Hayes (1981); Koo et al. (2023). Using this codebook, those annotators re-labeled each span of keystrokes with the corresponding label during the annotation process.
The annotators were fully informed about all the labels and had complete access to them when annotating each data point. The annotation process for all the labels is the same: First, they view through multiple consecutive data points and identify which high-level label occurs (e.g., Planning, Implementation, or Revision). Once annotators have identified the current high-level label, attempting to identify where it ends. Then, they decide on the low-level label within the high-level label (e.g., idea generation or organization under the Planning stage, etc.). Finally, they identify the interval for low-level labels and annotate data points in the interval with the identified low-level label. If a keystroke does not deliver any insight, then label it as an ‘artifact.’
We calculated inter-annotator agreement using the weighted F1 score in a multi-label, multi-class setting, which is suitable for our complex annotation schema involving multiple labels per instance. The weighted F1 score achieved was 0.71, indicating a high level of agreement between the annotators.
Appendix C More About the Taxonomy of Scholarly Writing Process
During the planning stage, the writers engage in a process of generating and organizing raw ideas, arguments, or content structures that were not introduced in the previous trajectory. Based on the plan, the writers implement their plan by drafting full sentences and paragraphs and structuring the contents tangibly. At the same time, the writers enter the revision stage by improving the quality of their implemented sentences and LaTeX objects in terms of linguistic styles, format, or information accuracy. Particularly, spans of keystrokes whose intentions involved any changes but did not change the meaning of original texts are classified as Revision. For those edits that show changes in the meaning, we considered them as Implementation. Furthermore, if an author repeatedly adds, removes, and revises text back and forth until a sentence is completed, we consider this process as part of text production. Any subsequent changes made to the sentence after it is finished are considered revision.
Appendix D More about ScholaWrite Evaluation
D.1 Training environments
BERT & RoBERTa
We fine-tuned BERT and RoBERTa with the following hyperparameter setups: (1) a learning rate of ; (2) training batch size per device of 8; (3) evaluation batch size per device of 8; (4) the number of training epochs of 10; and (5) a weight decay of . For each model, it took approximately 3.5 hours on one NVIDIA RTX A6000.
Llama
For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct111111https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library121212https://github.com/unslothai/unsloth.
For the intention prediction task (Sec. 5.1), here are the hyperparameter setups for the Llama models: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 5; (4) learning rate of ; and (5) AdamW 8-bit optimizer. Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset. For Llama-8B, it took approximately 8 hours on one RTX A5000.
For the ‘after’ text generation subtask from the iterative self-writing experiment (Sec. 5.2), we used the following hyperparameter setups for the fine-tuned Llama-8B-Instruct: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 10; (4) learning rate of ; and (5) AdamW 8-bit optimizer. Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset. Also, it took approximately 12 hours on one NVIDIA L40s.
GPT-4o
We used the GPT-4o-2024-08-06 version for the classification and iterative writing experiments. The iterative writing with 100 iterations took approximately 1 hour on each seed.
D.2 Details About Finetuning Process
Next Intention Prediction (Sec 5.1)
The fine-tuning prompt consisted of task instructions, a list of all intention labels with their definitions, the “before” text, and the corresponding human-annotated intention label. To achieve optimal performance while minimizing memory usage, we employed QLoRA Dettmers et al. (2024) to fine-tune all linear modules of a 4-bit quantized Llama. During fine-tuning, we utilized the ‘train_on_response_only’ function provided by the unsloth library, which masks the task instructions, intention label definitions, and “before” text with -100s. This ensures the model is trained exclusively on the response portion of the fine-tuning prompt (i.e., the predicted intention label), without being influenced by the instructional components of the input. The model was fine-tuned for one epoch with a batch size of 2, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
Iterative Self-Writing (Sec 5.2)
The structure of the “after” text differs slightly to help the model learn where and what edits to make. We used the diff_match_patch library to generate a word-level difference array between the “before” and “after” texts. Special tokens (<same>, </same>, <del>, </del>, <add>, </add>) were added to the tokenizer, and the difference array was converted into text wrapped with these tokens. For example, given a “before” text of ‘‘Bad dog’’ and an “after” text of ‘‘Good dog’’, the difference array would be [(-1, ‘Bad’), (1, ‘Good ’), (0, ‘dog’)]. This is converted into: <del>Bad </del><add>Good </add><same>dog</same>. This transformation was applied only to the “after” text, while the “before” text remained as plain LaTeX text.
The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text. For fine-tuning, we used QLoRA Dettmers et al. (2024) to optimize all linear modules of a 4-bit quantized model while maintaining a small memory footprint. Additionally, the embed_tokens and lm_head modules were set as trainable and saved in the final checkpoint. To focus training on the response portion (the “after” text), we used the train_on_response_only function, masking the task instructions, verbalizer, and “before” text with -100s. This ensures the model learns to generate the "after" text without being influenced by instructional input. The model was trained for one epoch with a batch size of 1, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
During iterative writing, we performed 100 iterations, treating the model’s output under one intention as a single iteration. If the intention predicted by the classification model (fine-tuned Llama3.1-8B-Instruct, as described in Sec 5.1) matched the current predicted intention, the model was prompted to edit the text again. In this case, the newly generated output was not treated as final output in the iteration, and the iteration did not proceed. We moved to the next iteration only when the intention prediction model generated a different intention label than the previous one.
For model setups, we created three pairs of models for each prediction and generation subtasks as follows:
LLama-8B-SW: LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Prediction) & LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Generation), independently
LLama-8B-Instruct: Vanilla LLama-8B-Instruct (Prediction) & Vanilla LLama-8B-Instruct (Generation), independently
GPT-4o: GPT-4o inference (Prediction) & GPT-4o inference (Generation)
D.3 Fine-tuning Prompt Templates
D.3.1 Prediction Prompt for Llama-8B-Zero models
“Here are all the possible writing intention labels:
Idea Generation: Formulate and record initial thoughts and concepts.
Idea Organization: Select the most useful materials and demarcate those generated ideas in a visually formatted way.
Section Planning: Initially create sections and sub-level structures.
Text Production: Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source.
Object Insertion: Insert visual claims of their arguments (e.g., figures, tables, equations, footnotes, itemized lists, etc.).
Cross-reference: Link different sections, figures, tables, or other elements within a document through referencing commands.
Citation Integration: Incorporate bibliographic references into a document and systematically link these references using citation commands.
Macro Insertion: Incorporate predefined commands or packages into a LaTeX document to alter its formatting.
Fluency: Fix grammatical or syntactic errors in the text or LaTeX commands.
Coherence: Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
Structural: Improve the flow of information by modifying the location of texts and objects.
Clarity: Improve the semantic relationships between texts to be more straightforward and concise.
Linguistic Style: Modify texts with the writer’s writing preferences regarding styles and word choices, etc.
Scientific Accuracy: Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims.
Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
Identify the most likely next writing intention of a graduate researcher when editing the following LaTex paper draft. Your output should only be a label from the list above.
## Input: before_text
## Output:
D.3.2 Generation Prompt for Llama-8B-Zero models
“You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
Below is the paper you have written so far. Please strictly follow the writing intention given above and insert, delete, or revise at appropriate places in the paper given below.
Your writing should relate to the paper given below. Do not generate text other than paper content. Do not describe the changes you are making or your reasoning.
## Input: before_text ”
D.3.3 Prediction Prompt for Llama-8B-SW models
“Here are all the possible writing intention labels:
Idea Generation: Formulate and record initial thoughts and concepts.
Idea Organization: Select the most useful materials and demarcate those generated ideas in a visually formatted way.
Section Planning: Initially create sections and sub-level structures.
Text Production: Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source.
Object Insertion: Insert visual claims of their arguments (e.g., figures, tables, equations, footnotes, itemized lists, etc.).
Cross-reference: Link different sections, figures, tables, or other elements within a document through referencing commands.
Citation Integration: Incorporate bibliographic references into a document and systematically link these references using citation commands.
Macro Insertion: Incorporate predefined commands or packages into a LaTeX document to alter its formatting.
Fluency: Fix grammatical or syntactic errors in the text or LaTeX commands.
Coherence: Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
Structural: Improve the flow of information by modifying the location of texts and objects.
Clarity: Improve the semantic relationships between texts to be more straightforward and concise.
Linguistic Style: Modify texts with the writer’s writing preferences regarding styles and word choices, etc.
Scientific Accuracy: Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims.
Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
Identify the most likely next writing intention of a graduate researcher when writing the following LaTex paper draft. Your output should only be a label from the list above.
## Input: before_text
## Output:
D.3.4 Generation Prompt for Llama-8B-SW models
“You are a computer science researcher with extensive experience in scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX.
You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
Below is the paper you have written so far. Given the paper information below and the corresponding scholarly writing intention, please revise or add to the text to fulfill this writing intention.
You may insert, delete, or revise text at appropriate places in the given paper.
Please provide a complete output. Do not generate text that is nonsensical or unrelated to the given paper information.
## Input: before_text
## Output:
D.3.5 Classification Prompt for GPT-4o
“You are a classifier that identify the most likely next writing intention. You will be given a list of all possible writing intention labels with definitions, and an in-progress LaTex paper draft written by a graduate student. Please strictly follow user’s instruction to identify the most likely next writing intention
Here is a list of all the possible writing intention labels with definitions:
Idea Generation: Formulate and record initial thoughts and concepts.
Idea Organization: Select the most useful materials and demarcate those generated ideas in a visually formatted way.
Section Planning: Initially create sections and sub-level structures.
Text Production: Translate their ideas into full languages, either from the writers’ language or borrowed sentences from an external source.
Object Insertion: Insert visual claims of their arguments (e.g., figures, tables, equations, footnotes, itemized lists, etc.).
Cross-reference: Link different sections, figures, tables, or other elements within a document through referencing commands.
Citation Integration: Incorporate bibliographic references into a document and systematically link these references using citation commands.
Macro Insertion: Incorporate predefined commands or packages into a LaTeX document to alter its formatting.
Fluency: Fix grammatical or syntactic errors in the text or LaTeX commands.
Coherence: Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
Structural: Improve the flow of information by modifying the location of texts and objects.
Clarity: Improve the semantic relationships between texts to be more straightforward and concise.
Linguistic Style: Modify texts with the writer’s writing preferences regarding styles and word choices, etc.
Scientific Accuracy: Update or correct scientific evidence (e.g., numbers, equations) for more accurate claims.
Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
Identify the most likely next writing intention of a graduate researcher when writing the following LaTex paper draft. Your output should only be a label from the list above.
Here is LaTeX paper draft:
## Input: before_text
## Output:
D.3.6 Generation Prompt for GPT-4o
“ You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
Your writing intention is to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
Below is the paper you have written so far. Please strictly follow the writing intention given above and insert, delete, or revise at appropriate places in the paper given below.
Your writing should relate to the paper given below. Do not generate text other than paper content. Do not describe the changes you are making or your reasoning. Do not include sidenotes. Your output should only be the paper draft in latex, without the “‘latex delimiters.
## Input: before_text
## Output:
D.4 Seed Documents for Iterative Self-Writing
We present the remaining three seed documents that we used for the iterative self-writing experiments, as shown in Listings LABEL:table:seed-entry-johnny to LABEL:table:seed-entry-latxa.

### v3

@@
 To fully understand writers’ cognitive thought process, one should fully decode the end-to-end writing data (from individual ideas to final manuscript) and understand their complex cognitive mechanisms in scholarly writing.
-We introduce ScholaWrite dataset, the first-of-its-kind keystroke logs of an end-to-end scholarly writing process for complete manuscripts, with thorough annotations of cognitive writing intentions behind each keystroke.
+We introduce ScholaWrite dataset, a first-of-its-kind keystroke corpus of an end-to-end scholarly writing process for complete manuscripts, with thorough annotations of cognitive writing intentions behind each keystroke.
 Our dataset includes LaTeX-based keystroke data from five preprints with nearly 62K total text changes and annotations across 4 months of paper writing.
-ScholaWrite shows promising usability and applications (e.g., iterative self-writing) for the future development of AI writing assistants for academic research, which necessitate complex methods beyond LLM prompting.
-Our experiments clearly demonstrated the importance of collection of end-to-end writing data, rather than the final manuscript, for the development of future writing assistants to support the cognitive thinking process of scientists.
-Our de-identified dataset, demo, and code repository are available on our project page111https://minnesotanlp.github.io/scholawrite/.
+ScholaWrite shows promising usability and applications (e.g., iterative self-writing), demonstrating the importance of collection of end-to-end writing data, rather than the final manuscript, for the development of future writing assistants to support the cognitive thinking process of scientists.
+Our de-identified data examples and code are available on our project page111https://minnesotanlp.github.io/scholawrite/.
 ScholaWrite: A Dataset of End-to-End Scholarly Writing Process
@@
 (2023); Wengelin and Johansson (2023), the specific cognitive patterns in scholarly writing remain understudied.
-Our work observes the end-to-end writing process of scholarly writing, inspired by the method of keystroke collections that have been a major methodology to observe individual writing processes Baaijen et al.
+Our work observes the end-to-end cognitive process of scholarly writing, inspired by the method of keystroke collections that have been a major methodology to observe individual writing processes Baaijen et al.
 (2012); Latif (2008); Leijten and Van Waes (2013) in cognitive science or a few recent works in NLP at a small scale Koo et al.
@@
 (2024).
-We believe our work (ScholaWrite) is the first attempt to present a keystroke corpus of scholarly writing with annotations of cognitive writing processes, which were collected over multiple months and produced by early-career researchers.
-Furthermore, we developed novel data collection and annotation systems using Chrome extension, which improves data accessibility to the public and enables real-time collection of LaTeX-based keystrokes from multiple researchers.
+We believe our work (ScholaWrite) is the first attempt to present a keystroke corpus of scholarly writing with annotations of cognitive writing intentions, which were collected over multiple months and produced by early-career researchers.
+Furthermore, we developed novel data collection and annotation systems, which improves data accessibility to the public and enables real-time collection of LaTeX-based keystrokes from multiple researchers.
 We also present a comprehensive taxonomy of cognitive writing processes specific to the scholarly writing domain.
@@
 We present ScholaWrite, a curated dataset of nearly 62K LaTeX-based keystrokes that were turned into publications in the computer science domain, annotated by experts in linguistics and computer science.
-We develop a taxonomy of scholarly writing intentions, providing an overall understanding of how scholars tend to develop writing manuscripts.
+We develop a taxonomy of cognitive process of scholarly writing, providing an overall understanding of how scholars tend to develop writing manuscripts.
 Experiment results showed that the Llama-8B model fine-tuned on the ScholaWrite dataset achieved the high linguistic quality of the final writing.
@@
 2.1 Cognitive Theory of Writing Process
-Research on human writing significantly shifted from focusing solely on the final written product to exploring the cognitive stages of various writing phases Diederich (1974); Perl (1979); Krapels (1990); MacArthur and Graham (2016).
-Building on this process-oriented approach, Flower and Hayes (1981)’s cognitive writing theory outlines three key cognitive sub-processes grounded in empirical evidence: (1) planning; (2) translating; and (3) reviewing.
-These processes are complex, interconnected, and non-linear, with writers moving fluidly between them.
-Our work expands this model to develop our own taxonomy of more granular cognitive patterns of human writing, specifically in scholarly communication.
+Research on human writing shifted from analyzing final written products to examining cognitive processes across various writing phases Diederich (1974); Krapels (1990); MacArthur and Graham (2016).
+Building on this process-oriented approach, Flower and Hayes (1981)’s cognitive writing theory outlines three key sub-processes: (1) planning, (2) translating, and (3) reviewing.
+These interconnected and non-linear stages of human writing inform our work, where we expand this model to a more granular taxonomy of cognitive writing patterns in scholarly communication.
 Koo et al.
-(2023) presented a taxonomy of the scholarly writing process based on keystroke collections of just a short research plan over a thirty-minute session.
-Inspired by this work, we extend their approach to a larger-scale collection of keystrokes over months, whose final written products turned into research publications.
-Reviewed by a linguistics expert, our comprehensive taxonomy is built on top of prior literature and encompasses end-to-end writing trajectories of the cognitive process in a scholarly domain.
+(2023) introduced a taxonomy of scholarly writing process based on keystroke data from short research plans written in 30-minute sessions.
+Expanding on this, we perform a larger-scale keystroke collection over months, whose final products culminated in research publications.
+Expert-reviewed and grounded in prior literature, our taxonomy encompasses end-to-end cognitive trajectories of scholarly writing.
 2.2 Keystroke Loggers for Scholarly Writing
-Keystroke logging tools such as Inputlog allow researchers to observe digital writing without interfering with the writer’s process Chan (2017); Johansson et al.
+Keystroke logging tools (e.g., Inputlog) allow researchers to observe digital writing without disrupting the writing process Chan (2017); Johansson et al.
 (2010); Leijten and Van Waes (2013); Lindgren and Sullivan (2019).
-However, they tend to operate in a closed ecosystem, i.e., only running in MS Word.
-This makes the current tools less accessible, especially for the scientific communities who use other typesetting such as LaTeX.
-Moreover, it is not well-suited for collecting data from extended writing sessions.
-Currently, such keystroke logging studies have rarely examined scholarly communication settings that use LaTeX, especially for scientific writing purposes.
-To address this gap, our work implemented novel several interconnected systems that securely collect keystrokes in real-time over months from Overleaf, a widely-used online LaTeX editor, store them without privacy issues, and retrieve them for annotating writing intentions at span level.
-Our system workflow provides a more natural environment for observing cognitive writing activities by multiple researchers over long periods and across multiple sections without interrupt their writing.
+However, they are often restricted to a closed ecosystem, such as MS Word, and thus less accessible for scientific communities who use LaTeX.
+Moreover, they are not well-suited for collecting data from extended writing sessions.
+Current studies rarely examined scholarly communication that uses LaTeX, particularly for scientific writing.
+To address this gap, we developed systems that securely collect real-time keystrokes over months from Overleaf, a widely-used online LaTeX editor, while ensuring privacy.
+The system workflow enables span-level annotation of writing intentions, offering a natural, uninterrupted environment for studying cognitive writing activities over long periods and across multiple sections.
 2.3 Datasets for Scholarly Writing
-Previous work primarily focused on constructing datasets for understanding the writing process of scholarly domains, which vary in their content and scale.
-Much of the publicly available datasets analyzed linguistic style changes or grammatical edits during the revision process Du et al.
+Previous work primarily focused on constructing datasets to analyze scholarly writing processes, which vary in content and scale.
+Publicly available datasets tend to track linguistic style changes or grammatical edits during revision Du et al.
 (2022b); Jiang et al.
@@
 (2019); Mita et al.
-(2022).
-Also, the datasets from D’Arcy et al.
+(2022), while others capture edits based on feedback and peer review D’Arcy et al.
 (2024); Jourdan et al.
 (2024); Kuznetsov et al.
-(2022) collected edits conditioned on feedback and peer review process, while Kobayashi et al.
+(2022) or citation generation Kobayashi et al.
 (2022); Narimatsu et al.
-(2021) constructed datasets for citation generations.
-Existing work is also limited to only specific sections of scientific papers (e.g., abstract, introduction) Du et al.
+(2021).
+Also, most focus on specific sections of papers (e.g., abstracts, introductions) Du et al.
 (2022b); Mita et al.
 (2022).
-In comparison, we construct a dataset that covers all cognitive phases of the writing process from its beginning to the final written product, through long periods of the data collection process.
-Furthermore, those existing corpora for scientific writing research are created from the comparison between multiple versions of full-length, final products retrieved from preprint databases Jiang et al.
+In contrast, our dataset covers all cognitive phases of writing, from inception to final manuscript, over an extended period.
+Furthermore, those existing corpora often compare multiple versions of final manuscripts from preprint databases Jiang et al.
 (2022); Du et al.
@@
 (2024); Jourdan et al.
+(2024), missing how those manuscripts evolved Jourdan et al.
 (2023).
-As pointed out by Jourdan et al.
-(2023), the comparison of finalized outputs does not present how those manuscripts evolved.
-Our work addresses this problem by building a corpus of keystrokes each of which represents the progress of the publications (See Figure 2).
+To address this, we build a keystroke-based corpus that captures the progression of publications in real-time (See Figure 2).
 3 Dataset Construction
-To preserve writers’ natural thought flow and collect authentic writing data, we developed a Chrome extension that records keystroke data in the background.
+To collect authentic writing data from writers’ natural thoughts, we developed a Chrome extension that records keystroke data in the background.
 After the study, an in-depth annotation process is conducted to interpret the intent behind each keystroke.
-For open research, we make all the code of our data collection and annotation framework publicly available: https://minnesotanlp.github.io/scholawrite/.
-We describe the detailed process of our data collection, annotation system design and taxonomy creation in the following sections.
+We describe the detailed process of our data collection, annotation system design, and taxonomy creation in the following sections.
 3.1 Participant Recruitment
-We recruited 10 graduate students222All of them currently attend a four-year university in the U.S.
-and are proficient in English.
-from the computer science department who were actively preparing academic manuscripts in Overleaf LaTeX editor333https://overleaf.com/ for submission to peer-reviewed conferences in AI and NLP.
+We recruited 10 graduate students222All participants attend a R1 university in US and are proficient in English.
+Appendix A.1 details the recruitment process.
+from the computer science department who were actively preparing academic manuscripts in Overleaf LaTeX editor for submission to peer-reviewed conferences in AI and NLP.
 The data is collected from November 2023 to February 2024, totaling up to 4 months.
@@
 3.2 Chrome Extension for Keystroke Collection in Overleaf
-We designed and implemented a Chrome extension (Figure 4(a)(a)), which enables the real-time collection of keystroke trajectories in the Overleaf platform.
-Participants can create their account credentials, and after logging into the system, the extension monitors the user’s keystrokes in the background silently, without disrupting the typical writing process.
-When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel.
-When each of these actions444Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page.
-occurs, the extension uses ‘diff_match_patch’ package555https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 3).
-Then, the extension will send the difference array along with other metadata (e.g., time stamp, author ID, file name, etc.) to the backend server.
-To prevent any issue of private data collection, the backend fetches only the IDs of the Overleaf projects that participants consented to share during the recruitment process and filters out participants’ keystroke data from any unauthorized projects666We used Google Sheet API to retrieve ID information that we collected during the recruitment process..
-For any Overleaf project that consists of multiple LaTex files, which is common in collaborative scholarly writing, we also collected all keystrokes from subfiles associated with the main LaTeX file.
-Our comprehensive data collection process captures the end-to-end writing processes of the participating authors across all parts of Overleaf projects.
-This approach ensures that our dataset reflects the full scope of scholarly writing including edits made in auxiliary files such as files of each section, appendix, bibliography, etc.
-All keystroke data stored in the database were processed for annotation and model training purposes.
-Here, raw keystrokes from the Chrome extension system are stored as arrays that represent the differences between subsequent keystrokes, along with accompanying metadata.
+We designed and implemented a Chrome extension (Appendix Figure 7), which enables the real-time collection of keystroke trajectories in the Overleaf platform.
+Participants can create their account credentials, and after logging into the system, the extension monitors the user’s keystrokes in the background silently without disrupting writing process.
+See Appendix A.2 for more description of the extension workflow.
 1st | Intention | Definition | An example action | Prop. |
-| Planning | Idea Generation | Formulate and record initial thoughts and concepts.
-| writing down keywords of a paragraph beforehand (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
+Planning | Idea Generation | Formulate and record initial thoughts and concepts. | writing keywords or notes (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
 Idea Organization | Select the most useful materials and demarcate those generated ideas in a visually formatted way.
@@
 3.3 Writing Intention Annotation
-To capture comprehensive writing behaviors, we designed a data collection process that involves running a Chrome extension in the background.
-However, keystroke data alone may be insufficient to fully understand the nuanced intentions behind each writing action.
-To bridge this gap, we implement an additional annotation step to decode the writer’s intent for each keystroke using a specialized annotation interface.
+To decode the nuanced intentions behind each writing action, we annotate each keystroke collected from the Chrome extension, using a specialized interface.
 This process is crucial for developing a detailed taxonomy of the end-to-end scholarly writing process, which serves as the basis for annotating the collected keystrokes.
-Further details about the annotation process can be found in Appendix B.
-To perform annotations, we first processed those raw keystroke entries by file name, type of writing actions (e.g., add, delete), an array of differences between two subsequent texts, and line number in an Overleaf editor.
+To perform annotations, we processed those raw keystroke entries by file name, type of writing actions, an array of differences between two subsequent texts, and line number in an Overleaf editor.
+Please see Appendix B for further annotation details.
 Annotation Interface
-We developed a novel interface for the keystroke annotation process.
-The interface offers various modes for visualizing keystroke collections within each Overleaf project: by time, LaTeX file, and author (Figure 4(a)(b)).
-In the right panel, annotators can interactively right-click to navigate the timeline of collected keystrokes in different colors (e.g., red for deletions and green for additions).
-They then identify a range of keystrokes whose underlying purpose or goal is similar when making edits during the writing process (e.g., generating outlines or correcting grammatical errors).
-They can manually enter the beginning and ending indices and annotate the span by selecting a pre-defined label of the relevant intention from the dropdown menu in the left panel.
-Furthermore, the interface shows the metadata information of each keystroke, along with the selected label.
+We developed a novel interface for keystroke annotation, enabling visualization by time, LaTeX file, and author (Appendix Figure 8).
+Annotators can navigate the keystroke timeline, identify a span of related edits (e.g., outlining or grammar correction), and assign a pre-defined intention label from the dropdown menu.
 Intention Taxonomy Construction
@@
 Following Pustejovsky et al.
-(2017)’s iterative annotate-and-refine framework, we performed an iterative open coding for the collected keystrokes from the two Overleaf projects.
-Using the annotation interface, each annotator reviewed the first few writing keystrokes and identified the high-level processes based on (Flower and Hayes, 1981)’s initial taxonomy of cognitive writing (planning, implementation, and revision).
+(2017), we performed an iterative open coding for the collected keystrokes from the two Overleaf projects.
+Using the annotation interface, each annotator reviewed the first few writing keystrokes and identified the high-level processes.
 Next, the annotators (1) determine the span of keystrokes that form a meaningful unit together (e.g., a clause or sentence) within the high-level process and (2) decide the underlying intention of edits.
@@
 Finally, a cognitive linguist participated in all discussions, refining the definition and scope of each writing intention.
-To ensure the validity of our taxonomy, we followed the two standard practices Nickerson et al.
+To ensure taxonomy validity, we followed two standard practices Nickerson et al.
 (2013); Kundisch et al.
-(2021): (1) All elements should be mutually exclusive and (2) the taxonomy should be collectively comprehensive.
-To achieve this, we flagged any ambiguous keystroke spans that could be interpreted as representing multiple intentions.
-Through iterative discussion, annotators refined these spans’ characteristics and improved each category’s decision boundaries.
-Once the codebook and taxonomy were finalized, we updated intention annotations777The list of those intentions is then updated in the dropdown bar of the annotation interface.
-for the selected two projects and applied the taxonomy to the remaining projects.
-The final inter-annotator agreement is using a weighted F1 score in a multi-label, multi-class setting for a randomly selected window of 1K keystrokes.
-Please refer to Appendix B for more details about the entire annotation procedures.
+(2021): (1) mutual exclusivity and (2) collective comprehensiveness.
+Ambiguous keystroke spans were flagged and refined through iterative discussions to sharpen category boundaries.
+After finalizing the codebook and taxonomy, we updated intention annotations333The updated list appears in annotation interface.
+for the two selected projects and applied the taxonomy to the rest.
+The final inter-annotator agreement was (weighted F1 score in a multi-label setting for a 1K-keystroke sample444Unlike Kappa, which assumes categorical agreement and struggles with multi-label classification, weighted F-1 scores better handle class imbalances and partial agreements.).
+See Appendix B for full annotation details.
 4 ScholaWrite: A Dataset of Cognitive Process of Scholarly Writing
-| Project | 1 | 2 | 3 | 4 | 5 |
-| # Authors | 1 (3) | 1 (4) | 1 (3) | 1 (4) | 9 (18) |
-| # keystrokes | 14,217 | 5,059 | 6,641 | 8,348 | 27,239 |
-| # Words added | 17,387 | 23,835 | 7,779 | 12,448 | 57,511 |
-| # Words deleted | 11,739 | 15,158 | 2,308 | 7,621 | 25,853 |
-| 1 | 2 | 3 | 4 | 5 | |
-| Idea Generation | 515 | 130 | 116 | 309 | 3255 |
-| Idea Organization | 0 | 45 | 25 | 9 | 231 |
-| Section Planning | 182 | 57 | 111 | 201 | 773 |
-| Text Production | 9267 | 2438 | 5109 | 4478 | 14031 |
-| Object Insertion | 583 | 383 | 62 | 486 | 1300 |
-| Cross-reference | 141 | 112 | 13 | 292 | 458 |
-| Citation Integration | 75 | 151 | 69 | 127 | 245 |
-| Macro Insertion | 16 | 7 | 51 | 29 | 33 |
-| Linguistic Style | 233 | 75 | 42 | 201 | 411 |
-| Coherence | 422 | 242 | 126 | 193 | 1021 |
-| Clarity | 1249 | 645 | 721 | 1180 | 3301 |
-| Scientific Accuracy | 307 | 15 | 2 | 24 | 95 |
-| Structural | 359 | 506 | 105 | 257 | 1042 |
-| Fluency | 116 | 90 | 46 | 135 | 476 |
-| Visual Formatting | 752 | 163 | 43 | 427 | 567 |
 Summary
-Our post-processed dataset, ScholaWrite, contains end-to-end writing trajectories with intentions annotated, for five Overleaf projects whose final product turned into arXiv preprints.
+Our post-processed dataset, ScholaWrite, contains end-to-end writing trajectories with annotated intentions, for five Overleaf projects whose final product turned into arXiv preprints.
 The dataset consists of 61,504 arrays of keystroke changes with fine-grained annotations of writing intentions.
-We publicly release our post-processed data in Huggingface data cards.888https://huggingface.co/datasets/minnesotanlp/scholawrite Table 3 shows the overall statistics of the dataset.
+Appendix Table 6 shows the overall statistics of the dataset.
 Writing Intention Distributions
-The writing intention distributions is illustrated in Figure 5(a).
+The writing intention distributions is illustrated in Figure 3(c).
 During the planning stage, idea generation is the most frequent intention.
@@
 Since we defined the boundary between implementation and revision as the point where meaning changes occur, the high frequency of text production suggests that authors often continue revising their texts by introducing new topics into the documents.
-Table 3 shows the label distribution per project.
+Appendix Table 8 shows the label distribution per project.
+Label | Subsequent label | Probability | |
+Idea Generation | Text Production | 0.52 | |
+Idea Organization | Idea Generation | 0.34 | |
+Citation Integration | Text Production | 0.37 | |
+Cross-reference | Text Production | 0.36 | |
+Clarity | Text Production | 0.35 | |
+Coherence | Text Production | 0.34 | |
+Scientific Accuracy | Text Production | 0.34 |
 Flows Between Writing Intentions
@@
 Many intentions proceed to text production, where the idea generation has the highest probability of feeding into text production.
-Also, the reflexive relationship between text production and clarity suggests a constant loop of producing new texts and refining them, ensuring the precision and reasonability of their claims.
-| Label | Subsequent label | Probability | |
-| Idea Generation | Text Production | 0.52 | |
-| Idea Organization | Idea Generation | 0.34 | |
-| Section Planning | Text Production | 0.33 | |
-| Text Production | Clarity | 0.20 | |
-| Object Insertion | Text Production | 0.32 | |
-| Citation Integration | Text Production | 0.37 | |
-| Cross-reference | Text Production | 0.36 | |
-| Macro Insertion | Idea Generation | 0.29 | |
-| Fluency | Text Production | 0.30 | |
-| Coherence | Text Production | 0.34 | |
-| Clarity | Text Production | 0.35 | |
-| Structural | Text Production | 0.27 | |
-| Linguistic Style | Text Production | 0.29 | |
-| Scientific Accuracy | Text Production | 0.34 | |
-| Visual Formatting | Text Production | 0.25 |
+Figure 3(a) also shows a similar pattern of flows between text productions and other intentions.
+Also, in Table 2 the reflexive relationship between text production and clarity suggests a constant loop of producing new texts and refining them, ensuring the precision and reasonability of their claims.
+Please see Appendix Figure 13 for the intention flows of each of the five projects.
 Time-series Distributions of Writing Intentions
 We calculated the average Wasserstein distance between each intention distribution and uniform distribution to assess how evenly intentions occur throughout the writing process.
-As shown in Figure 5(b), text production had the smallest distance, while macro insertion had the largest.
-This suggests that text production is spread throughout the writing, whereas macro insertion is concentrated in specific phases.
-A similar pattern appears in Figure 5(c), where text production is evenly distributed over time, and macro insertion peaks in the early and final stages of the writing process.
-Figures 6(a) and 6(b) demonstrate a similar pattern that text production of the implementation stage spreads evenly over time.
+As shown in Figure 3(d), text production had the smallest distance, while scientific accuracy had the largest.
+This suggests that text production is spread throughout the writing, whereas scientific accuracy is concentrated in specific phases.
+A similar pattern appears in Figure 3(e) , where text production is evenly distributed over time, and scientific accuracy peaks in the middle-to-late of the writing process.
+Figure 3(b) demonstrates a similar pattern that text production of the implementation stage spreads evenly over time.
 Types of clarity and structural revisions tend to occur after the medium steps of writing.
-Please refer to Figures 9 to 12 in the Appendix for each of the five Overleaf projects.
-5 Applications for Writing Assistance with ScholaWrite
-We believe ScholaWrite can serve as a valuable resource for training language models or enhancing the cognitive writing process in future writing assistants for scholarly writing.
-To evaluate its usability, we performed several experiments where existing LLMs were trained to mimic the complex, non-linear scientific writing processes of human scholars.
+Please see Appendix Figures 9 to 12 for each of the five projects.
+5 Applications for Writing Assistance
+We envision ScholaWrite as a valuable resource for training language models and improving future writing assistants for scholarly writing.
+To evaluate its usability, we conducted experiments training LLMs to mimic the complex, non-linear writing processes of human scholars.
 Specifically, we aimed to showcase the capabilities of LLMs trained on ScholaWrite in two scenarios:
-(1) Predicting an author’s next writing intention (Section 5.1)
-: The task is crucial for writing assistants to accurately assess the writer’s current status in context and predict the correct writing intention.
+(1) Predicting an author’s next writing intention (§5.1): The task is crucial for writing assistants to accurately assess the writer’s current status in context and predict the correct writing intention.
 This enables them to offer cognitively-appropriate writing suggestions that align with the writer’s needs.
-(2) Iteratively generating scholarly writing actions from scratch (called Iterative Self-Writing), mirroring the human writing process (Section 5.2)
-: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
-As illustrated in Figure 7, the first task (represented by the inner box labeled "Prediction") takes the "before" text from a keystroke pair and context prompt, and predicts the writing intention to apply for the subsequent actions.
-Once the intention is predicted, the second task generates the “after” text based on the predicted intention and given prompt (represented by the outer box labeled “Generation”).
-This process repeats iteratively until no further changes are made or a maximum number of iterations, such as 100, is reached.
-We randomly split our dataset into train (80%) and test (20%) sets.
-For each intention label in the test set, we randomly select up to 300 keystroke entries, due to budget constraints.
-Note that the same train and test sets are applied to all models across all experiments.
-See Appendix D for a full description of the model training process.
+(2) Iteratively generating scholarly writing actions from scratch (§5.2) (called Iterative Self-Writing), mirroring the human writing process: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
 5.1 Predicting Next Writing Intention
 Setup & Metrics
+This task (represented by the inner box “Prediction” in Figure 4) takes the “before-text” from a keystroke pair (i.e., Listing LABEL:table:single-entry) and context prompt, and predicts the writing intention to apply for the subsequent actions.
 We use BERT Devlin et al.
@@
 For comparison, we also run GPT-4o OpenAI (2024) on the test set.
-The fine-tuning prompt included all possible labels with definitions, task instructions, the “before” text chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”.
-Differences in prompts were limited to only task instructions (see Appendix D.3 for prompt details).
-To evaluate model performance, we used a weighted f-1 score999Weighted F-1 was chosen to address skewed label distribution (as shown in Figure 5(a) and Table 1)., comparing predictions to gold intention labels on the test set.
+To evaluate model performance, we used a weighted F-1 score555Weighted F-1 was chosen to address skewed label distribution (as shown in Figure 3(c) and Table 1)..
+See Appendix E.2 for details.
 BERT | RoBERTa | Ll ama-8B | GPT-4o | |
@@
 + SW | 0.64 | 0.64 | 0.13 | - |
+Iter. | Llama-8B-SW | Llama-8B-Zero | GPT-4o |
+[…]
+[…]
+[…]
+[…]
 Results
 Table 3 presents the weighted F1 scores for predicting writing intentions across baselines and fine-tuned models.
-Regardless of the intricate nature of the task itself101010Each model is asked to predict the next intention by only giving the ‘before’ text.
-Human annotators, however, label the data by looking through multiple consecutive ‘before and after’ edits to determine the current intention.
-In addition, the next intention chosen by the author does not necessarily mean that it is the only correct intention., all models finetuned on ScholaWrite show an improved performance compared to their baselines.
+Regardless of the intricate nature of the task itself666Each model predicts the next intention using only the ”before” text, while human annotators consider multiple ”before and after” edits.
+Moreover, the chosen next intention is not necessarily the only correct one., all models finetuned on ScholaWrite show an improved performance compared to their baselines.
 BERT and RoBERTa achieved the most improvement, while LLama-8B-Instruct showed a modest improvement after fine-tuning.
-As detailed in Appendix D, more training epochs and encoder-decoder architectures of BERT variants are assumed to be the reason for significant improvement compared to LLMs.
+As detailed in Appendix E, more training epochs and encoder-decoder architectures of BERT variants are assumed to be the reason for significant improvement compared to LLMs.
 This aligns with findings from Grasso and Locci (2024); Yu et al.
@@
 Setups
-Iterative self-writing involves two subtasks (Figure 7): (1) next intention prediction and (2) "after-text" generation.
-For intention prediction, we use fine-tuned Llama3.1-8B-Instruct on the ScholaWrite training set (“Llama-8B-SW-pred”), compared to vanilla models (Llama-8B-Instruct, and GPT-4o) from Section 5.1 as baselines.
-Please refer to Figure 13 in Appendix D for the detailed pipeline.
-For “after-text” generation, another vanilla Llama3.1-8B-Instruct is fine-tuned on the ScholaWrite training set (“Llama-8B-SW-gen”), with baselines including vanilla Llama3.1-8B-Instruct and GPT-4o.
-The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text.
-The model’s task is to generate “after-text” given the verbalizer and “before” text.
-To maintain consistency across outputs, prompts were adjusted for each model, with differences mainly in task instructions (see Appendix D.3 for the prompt templates).
-Seed documents were derived from four award-winning NLP papers spanning different topics Zeng et al.
+During iterative self-writing (Figure 4), a fine-tuned model processes LaTeX-formatted seed document (as “before-text”) with a context prompt to predict the next intention, then revises the text (“after-text”) accordingly given prompt.
+The revised document then serves as the new seed for the next iteration.
+This process repeats until a set iteration limit (e.g., 100) is reached.
+All models use the same train (80%)-test (20%) split across experiments.
+See Appendix E and Figure 14 for training details.
+We fine-tune Llama3.1-8B-Instruct (Llama-8B-SW) and compare it to vanilla Llama-8B-Instruct (Llama-8B-Zero) and GPT-4o.
+Also, seed documents were derived from LaTeX-formatted abstracts of four award-winning NLP papers on diverse topics Zeng et al.
 (2024); Lu et al.
@@
 (2022a); Etxaniz et al.
-(2024).
-Listing LABEL:table:seed-entry presents the LaTeX-formatted abstract of Du et al.
-(2022a) as an example seed text for initiating the iterative self-writing experiment.
+(2024) (Appendix LABEL:table:seed-entry-johnny-LABEL:table:seed-entry-latxa).
+Metrics
+We evaluated lexical diversity (unique-to-total token ratio), topic consistency (cosine similarity between seed and final output), and intention coverage (unique writing intentions used over 100 iterations).
+For human evaluation, three native English speakers with LaTeX expertise assessed outputs from Llama-8B-Zero and Llama-8B-SW on accuracy (alignment with predicted intention), alignment (similarity to human writing), fluency(grammatical correctness), coherence(logical structure), and relevance (connection to the seed document) - Refer to Appendix LABEL:sec:appendix:human-eval.
+Accuracy was judged per iteration, while other metrics used pairwise comparisons.
+Inter-annotator agreement (IAA)777The IAA scores are 0.84 (SW) and 0.76 (Zero) for accuracy, all 100% for alignment, fluency, coherence, and 49.8% (SW) and 100% (Zero) for relevance.
+was measured using Krippendorff’s alpha for accuracy and percentage agreement for others.
+Metrics | Model | Seed 1 | Seed 2 | Seed 3 | Seed 4 |
+Accuracy | SW | 21.0 | 10.3 | 18.3 | 15.3 |
+Zero | 35.7 | 29.7 | 45.3 | 43.3 | |
+Alignment | SW | 0 | 0 | 0 | 0 |
+Zero | 3 | 3 | 3 | 3 | |
+Fluency | SW | 0 | 0 | 0 | 0 |
+Zero | 3 | 3 | 3 | 3 | |
+Coherence | SW | 0 | 0 | 0 | 0 |
+Zero | 3 | 3 | 3 | 3 | |
+Relevance | SW | 1 | 3 | 2 | 1 |
+Zero | 3 | 3 | 3 | 3 |
+Results
+Figure 5 shows that Llama-8B-SW consistently produced the most lexically diverse words, generated the most semantically aligned topics (Seeds 1 & 2), and covered the most writing intentions (except Seed 3).
+These results underscore the value of ScholaWrite in improving scholarly writing quality generated by language models.
+However, our human evaluation (Table 5) revealed that Llama-8B-SW generated less human-like writing, in terms of fluency and logical claims.
+It also struggled with generating texts aligned with the predicted intentions.
+See Appendix Tables LABEL:table:human-eval-seed1 to LABEL:table:human-eval-seed4 for more details.
+Despite the weaknesses, Llama-8B-SW still produced more relevant content (Seed 2), which aligns with topic consistency trends in Figure 5, highlighting the usefulness of ScholaWrite dataset in certain contexts.
+Moreover, Llama-8B-SW exhibited the most human-like writing activity patterns over time (Figure LABEL:fig:writing-step-intention-all-model), which frequently switches between implementation and revision and covers all three high-level processes.
+Llama-8B-Zero and GPT-4o tend to remain in a single high-level stage throughout all 100 iterations of self-writing.
+Compared to Appendix Figure 12, which depicts frequent transitions across all three stages in an early draft (e.g., the first 100 steps), Llama-8B-SW most closely replicates human writing behaviors in iterative writing tasks.
+These findings reinforce the potential of ScholaWrite in helping LLMs emulate human scholarly writing processes.
 6 Conclusion and Future Work
-We introduce ScholaWrite, the first-of-its-kind dataset that captures the end-to-end scholarly writing process, consisting of nearly 62K LaTeX keystrokes collected via our custom-built Chrome extension.
+We present ScholaWrite, a first-of-its-kind dataset capturing the end-to-end cognitive process of scholarly writing, comprising nearly 62K LaTeX keystrokes collected via our custom-built Chrome extension.
 This dataset, sourced from ten graduate students with varying levels of scientific writing expertise, is further enriched with expert annotations based on a novel taxonomy of cognitive writing intentions inspired by Flower and Hayes (1981); Koo et al.
 (2023).
-Our study demonstrates the value of ScholaWrite as a resource for advancing the cognitive capabilities of large language models (LLMs) and developing cognitively-aligned writing assistants, enabling them to emulate the complex, non-linear, and intention-driven nature of human writing.
+Through several experiments, ScholaWrite shows its value for advancing the cognitive capabilities of LLMs and developing cognitively-aligned writing assistants, enabling them to mimic the complex, non-linear, and intention-driven nature of human writing.
 Future work includes expanding the dataset to diverse academic fields, authors, and collaborative projects, thus enabling models to generalize beyond fact recall to emulate human decision-making and reasoning in more realistic academic environments.
@@
 Importantly, no LLMs were used during any stage of the study, except for grammatical error correction in this manuscript.
-Acknowledgements
-This work was supported by the research gift from Grammarly.
-We thank Anna Martin-Boyle and Ryan Koo for their insights and the data collection tool based on their earlier version of this paper.
-We thank the participants of our data collection for their valuable time, effort, and contributions, which were essential to the success of this research.
-We also thank Minnesota NLP group members and anonymous reviewers for providing us with valuable feedback and comments on the paper draft.
 Appendix A More About Data Collection Process
@@
 A.2 Technical Details of System Implementations
+When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel888To prevent privacy concerns, the extension filters out keystroke data from any unauthorized Overleaf projects.
+Please see Appendix A.2 for more details..
+When each of these actions999Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page.
+occurs, the extension uses ‘diff_match_patch’ package101010https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 6).
+Then, the extension will send the array along with metadata (e.g., time stamp, author ID, etc.) to the backend server.
+For any Overleaf project that consists of multiple LaTeX files, we also collected all keystrokes from subfiles associated with the main LaTeX file.
+Our comprehensive data collection process captures the end-to-end writing processes of the participating authors across all parts of Overleaf projects.
+This approach ensures that our dataset reflects the full scope of scholarly writing including edits made in auxiliary files such as files of each section, appendix, bibliography, etc.
 We explain the technical implementation details of the two systems for the data collection process.
@@
 All data for the annotation interface was retrieved from the MongoDB database used in the Chrome extension system.
+Privacy Concerns
+To prevent any issue of private data collection, we designed the backend of our Chrome extension to fetch only the IDs of the Overleaf projects that participants consented to share during the recruitment process and filter out participants’ keystroke data from any unauthorized projects.
+We used Google Sheet API to retrieve ID information that we collected during the recruitment process.
 A.3 Data Post-Processing
@@
 Any subsequent changes made to the sentence after it is finished are considered revision.
-Appendix D More about ScholaWrite Evaluation
-D.1 Training environments
+Table 1 presents the comprehensive, complete definitions of each intention of end-to-end scholarly writing process, identified from ScholaWrite dataset.
+Appendix D ScholaWrite dataset statistics
+Table 6 presents the overall statistics of our ScholaWrite dataset. Table 8 shows the distribution of intention labels per Overleaf project from ScholaWrite.
+Project | 1 | 2 | 3 | 4 | 5 |
+# Authors | 1 (3) | 1 (4) | 1 (3) | 1 (4) | 9 (18) |
+# keystrokes | 14,217 | 5,059 | 6,641 | 8,348 | 27,239 |
+# Words added | 17,387 | 23,835 | 7,779 | 12,448 | 57,511 |
+# Words deleted | 11,739 | 15,158 | 2,308 | 7,621 | 25,853 |
+Label | Subsequent label | Probability | |
+Idea Generation | Text Production | 0.52 | |
+Idea Organization | Idea Generation | 0.34 | |
+Section Planning | Text Production | 0.33 | |
+Text Production | Clarity | 0.20 | |
+Object Insertion | Text Production | 0.32 | |
+Citation Integration | Text Production | 0.37 | |
+Cross-reference | Text Production | 0.36 | |
+Macro Insertion | Idea Generation | 0.29 | |
+Fluency | Text Production | 0.30 | |
+Coherence | Text Production | 0.34 | |
+Clarity | Text Production | 0.35 | |
+Structural | Text Production | 0.27 | |
+Linguistic Style | Text Production | 0.29 | |
+Scientific Accuracy | Text Production | 0.34 | |
+Visual Formatting | Text Production | 0.25 |
+1 | 2 | 3 | 4 | 5 | |
+Idea Generation | 515 | 130 | 116 | 309 | 3255 |
+Idea Organization | 0 | 45 | 25 | 9 | 231 |
+Section Planning | 182 | 57 | 111 | 201 | 773 |
+Text Production | 9267 | 2438 | 5109 | 4478 | 14031 |
+Object Insertion | 583 | 383 | 62 | 486 | 1300 |
+Cross-reference | 141 | 112 | 13 | 292 | 458 |
+Citation Integration | 75 | 151 | 69 | 127 | 245 |
+Macro Insertion | 16 | 7 | 51 | 29 | 33 |
+Linguistic Style | 233 | 75 | 42 | 201 | 411 |
+Coherence | 422 | 242 | 126 | 193 | 1021 |
+Clarity | 1249 | 645 | 721 | 1180 | 3301 |
+Scientific Accuracy | 307 | 15 | 2 | 24 | 95 |
+Structural | 359 | 506 | 105 | 257 | 1042 |
+Fluency | 116 | 90 | 46 | 135 | 476 |
+Visual Formatting | 752 | 163 | 43 | 427 | 567 |
+Figures 9 to 12 show several characteristics of human writing process, analyzed from ScholaWrite dataset: (1) Figure 9 - the average Wasserstein distance between each intention distribution and uniform distribution; (2) Figure 10 - distribution of labels over time; (3) Figure 11 for high-level intention distribution over time; and (4) Figure 12 for intention-wise writing activity distribution over time.
+Appendix E More about ScholaWrite Evaluation
+E.1 Training environments
 BERT & RoBERTa
@@
 We used the GPT-4o-2024-08-06 version for the classification and iterative writing experiments. The iterative writing with 100 iterations took approximately 1 hour on each seed.
-D.2 Details About Finetuning Process
+E.2 Details About Finetuning Process
 Next Intention Prediction (Sec 5.1)
-The fine-tuning prompt consisted of task instructions, a list of all intention labels with their definitions, the “before” text, and the corresponding human-annotated intention label.
+The fine-tuning prompt included all possible labels with definitions, task instructions, the “before-text” chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”.
+Differences in prompts were limited to only task instructions (see §E.3 for prompt details).
 To achieve optimal performance while minimizing memory usage, we employed QLoRA Dettmers et al.
@@
 This transformation was applied only to the “after” text, while the “before” text remained as plain LaTeX text.
+For finetuning, we randomly split the ScholaWrite dataset into training (80%) and testing (20%) sets.
+From each intention label in the test set, we sample up to 300 keystroke entries due to budget constraints.
+For intention prediction, we fine-tune Llama3.1-8B-Instruct on ScholaWrite training set (Llama-8B-SW-pred”) and compare it to baseline models (Llama-8B-Instruct and GPT-4o) from Section 5.1.
+For “after-text” generation, we fine-tune another Llama3.1-8B-Instruct model (“Llama-8B-SW-gen”) using the same dataset, with Llama3.1-8B-Instruct and GPT-4o as baselines.
+The fine-tuning prompt includes task instructions, a verbalizer from human-annotated labels, and “before-text.” While prompts were standardized, task instructions varied by model.
 The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text.
@@
 GPT-4o: GPT-4o inference (Prediction) & GPT-4o inference (Generation)
-D.3 Fine-tuning Prompt Templates
-D.3.1 Prediction Prompt for Llama-8B-Zero models
+Also, due to budget constraints, models had different revision strategies.
+Llama-8B-SW-* and Llama-8B-Instruct continued revision until the next predicted intention changed.
+GPT-4o, however, moved to the next iteration regardless.
+We refer to the fine-tuned Llama-8B model as Llama-8B-ScholaWrite (or Llama-8B-SW) and the vanilla model as Llama-8B-Zero.
+E.3 Prompt Templates
+E.3.1 Prediction Prompt for Llama-8B-Zero models
 “Here are all the possible writing intention labels:
@@
 ## Output:
-D.3.2 Generation Prompt for Llama-8B-Zero models
+E.3.2 Generation Prompt for Llama-8B-Zero models
 “You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
@@
 ## Input: before_text ”
-D.3.3 Prediction Prompt for Llama-8B-SW models
+E.3.3 Prediction Prompt for Llama-8B-SW models
 “Here are all the possible writing intention labels:
@@
 ## Output:
-D.3.4 Generation Prompt for Llama-8B-SW models
+E.3.4 Generation Prompt for Llama-8B-Zero models
 “You are a computer science researcher with extensive experience in scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX.
@@
 Please provide a complete output. Do not generate text that is nonsensical or unrelated to the given paper information.
-## Input: before_text
-## Output:
-D.3.5 Classification Prompt for GPT-4o
+## Input: before_text ”
+E.3.5 Classification Prompt for GPT-4o
 “You are a classifier that identify the most likely next writing intention.
 You will be given a list of all possible writing intention labels with definitions, and an in-progress LaTex paper draft written by a graduate student.
-Please strictly follow user’s instruction to identify the most likely next writing intention
-Here is a list of all the possible writing intention labels with definitions:
+Please strictly follow user’s instruction to identify the most likely next writing intention.
+Here are the verbalizers of all the possible writing intention labels:
 Idea Generation: Formulate and record initial thoughts and concepts.
@@
 Visual Formatting: Modify the stylistic formatting of texts, objects, and citations.
-Identify the most likely next writing intention of a graduate researcher when writing the following LaTex paper draft. Your output should only be a label from the list above.
-Here is LaTeX paper draft:
+Identify the most likely next writing intention of a graduate researcher when editing the following LaTex paper draft. Your output should only be a label from the list above.
 ## Input: before_text
 ## Output:
-D.3.6 Generation Prompt for GPT-4o
+E.3.6 Generation Prompt for GPT-4o
 “ You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
@@
 ## Output:
-D.4 Seed Documents for Iterative Self-Writing
+E.4 Seed Documents for Iterative Self-Writing
 We present the remaining three seed documents that we used for the iterative self-writing experiments, as shown in Listings LABEL:table:seed-entry-johnny to LABEL:table:seed-entry-latxa.

### v4

@@
 Abstract
-Writing is a cognitively demanding task involving continuous decision-making, heavy use of working memory, and frequent switching between multiple activities.
-Scholarly writing is particularly complex as it requires authors to coordinate many pieces of multiform knowledge.
-To fully understand writers’ cognitive thought process, one should fully decode the end-to-end writing data (from individual ideas to final manuscript) and understand their complex cognitive mechanisms in scholarly writing.
-We introduce ScholaWrite dataset, a first-of-its-kind keystroke corpus of an end-to-end scholarly writing process for complete manuscripts, with thorough annotations of cognitive writing intentions behind each keystroke.
-Our dataset includes LaTeX-based keystroke data from five preprints with nearly 62K total text changes and annotations across 4 months of paper writing.
-ScholaWrite shows promising usability and applications (e.g., iterative self-writing), demonstrating the importance of collection of end-to-end writing data, rather than the final manuscript, for the development of future writing assistants to support the cognitive thinking process of scientists.
-Our de-identified data examples and code are available on our project page111https://minnesotanlp.github.io/scholawrite/.
+Writing is a cognitively demanding activity that requires constant decision-making, heavy reliance on working memory, and frequent shifts between tasks of different goals.
+To build writing assistants that truly align with writers’ cognition, we must capture and decode the complete thought process behind how writers transform ideas into final texts.
+We present ScholaWrite, the first dataset of end-to-end scholarly writing, tracing the multi-month journey from initial drafts to final manuscripts.
+We contribute three key advances: (1) a Chrome extension that unobtrusively records keystrokes on Overleaf, enabling the collection of realistic, in-situ writing data; (2) a novel corpus of full scholarly manuscripts, enriched with fine-grained annotations of cognitive writing intentions.
+The dataset includes LaTeX-based edits from five computer science preprints, capturing nearly 62K text changes over four months; and (3) analyses and insights into the micro-dynamics of scholarly writing, highlighting gaps between human writing processes and the current capabilities of large language models (LLMs) in providing meaningful assistance.
+ScholaWrite underscores the value of capturing end-to-end writing data to develop future writing assistants that support, not replace, the cognitive work of scientists.
 ScholaWrite: A Dataset of End-to-End Scholarly Writing Process
-Linghe Wang††thanks: Equal contribution. Minhwa Lee Ross Volkov Luan Tuyen Chau Dongyeop Kang University of Minnesota {wang9257,lee03533,volko032,chau0139,dongyeop}@umn.edu
+Khanh Chi Le††thanks: Equal Contribution, Linghe Wang∗, Minhwa Lee∗, Ross Volkov, Luan Tuyen Chau, Dongyeop Kang University of Minnesota {le000422, wang9257, lee03533, volko032, chau0139, dongyeop}@umn.edu
 1 Introduction
-Scholarly writing requires researchers to produce texts that concisely and precisely present novel findings and to follow a systematized structure and style by their targeted venue Jourdan et al.
-(2023); Kallestinova (2011); Bourekkache (2022).
-To address this need, researchers have leveraged large language models (LLMs) to develop intelligent support systems in several writing tasks, such as revision process Du et al.
-(2022b); Kim et al.
-(2022) or feedback generation Liang et al.
-(2024); Lu et al.
-(2024a).
-However, a critical question arises: Are LLMs able to truly understand and generate writing as human scientists do?
-LLMs are generally trained to progress autoregressively, generating text from left to right.
-In contrast, human writing typically involves multiple iterations of complex and non-linear cognitive actions to refine the main message Flower and Hayes (1981); Koo et al.
-(2023), as illustrated in Figure 1.
-This fundamental cognitive gap between human writing and language models necessitates a deeper understanding of human writing processes and the development of cognitively-aligned writing assistance tools, moving beyond basic auto-completion.
-Although much literature has examined the cognitive processes used in general writing tasks Lindgren and Sullivan (2003); Hayes and Berninger (2014); Zhu et al.
-(2023); Wengelin and Johansson (2023), the specific cognitive patterns in scholarly writing remain understudied.
-Our work observes the end-to-end cognitive process of scholarly writing, inspired by the method of keystroke collections that have been a major methodology to observe individual writing processes Baaijen et al.
-(2012); Latif (2008); Leijten and Van Waes (2013) in cognitive science or a few recent works in NLP at a small scale Koo et al.
-(2023); Velentzas et al.
-(2024).
-We believe our work (ScholaWrite) is the first attempt to present a keystroke corpus of scholarly writing with annotations of cognitive writing intentions, which were collected over multiple months and produced by early-career researchers.
-Furthermore, we developed novel data collection and annotation systems, which improves data accessibility to the public and enables real-time collection of LaTeX-based keystrokes from multiple researchers.
-We also present a comprehensive taxonomy of cognitive writing processes specific to the scholarly writing domain.
-Throughout several experiments with LLMs, ScholaWrite shows promising usability as a resource for enabling LLMs to understand the human cognitive process of scholarly writing and provide writing suggestions that are aware of writers’ cognitive behaviors.
-Contributions
-We present ScholaWrite, a curated dataset of nearly 62K LaTeX-based keystrokes that were turned into publications in the computer science domain, annotated by experts in linguistics and computer science.
-We develop a taxonomy of cognitive process of scholarly writing, providing an overall understanding of how scholars tend to develop writing manuscripts.
-Experiment results showed that the Llama-8B model fine-tuned on the ScholaWrite dataset achieved the high linguistic quality of the final writing.
+Scientific writing is one of the most cognitively demanding forms of human communication.
+Researchers must articulate novel ideas with precision and structure, iteratively refining arguments, evidence, and phrasing over time (Jourdan et al., 2023; Kallestinova, 2011; Bourekkache, 2022).
+Unlike linear text generation, scholarly writing is a complex cognitive process, writers continually plan, draft, and revise across multiple stages and intentions (see Fig.
+1).
+Yet, despite the growing integration of large language models (LLMs) into research workflows, our understanding of how human scientists actually write remains limited.
+Without such understanding, it is difficult to build writing assistants that genuinely align with the cognitive processes of their users.
+Recent efforts in writing assistance have leveraged LLMs for tasks such as revision or feedback generation (Du et al., 2022b; Liang et al., 2024).
+However, these models largely operate autoregressively, producing text from left to right, while human writing unfolds through recursive, non-linear cycles of ideation, organization, and refinement (Flower and Hayes, 1981).
+This fundamental cognitive gap raises a key question: how can we empirically capture and model the end-to-end scholarly writing process so that future LLMs can support, rather than substitute, human cognition?
+Existing studies have examined writing processes through revision logs or small-scale keystroke analyses (Leijten and Van Waes, 2013; Koo et al., 2023), but most focus on short writing tasks or version-level comparisons of finalized drafts (Du et al., 2022b; Jiang et al., 2022; D’Arcy et al., 2024).
+Such approaches miss the fine-grained temporal and cognitive dynamics that drive real-world scholarly writing over extended periods.
+To bridge this gap, we present ScholaWrite, the first dataset capturing the end-to-end scholarly writing process, from initial drafts to final manuscripts, through in-situ keystroke logging and cognitive annotation.
+Our work contributes three main advances:
+(1) A reusable tool and taxonomy for cognitive writing capture.
+We develop a Chrome extension that unobtrusively records keystrokes within Overleaf, the dominant platform for collaborative scientific writing.
+Paired with a taxonomy of fifteen cognitive writing intentions, this tool enables large-scale, domain-specific capture of authentic writing behaviors (§3).
+(2) The ScholaWrite dataset.
+Using this system, we collect over 61K keystroke-based text changes from five Computer Science preprints written by graduate researchers over four months.
+Each keystroke is annotated with its underlying cognitive intention (e.g., idea generation, clarity improvement), producing a richly structured corpus for modeling writing cognition (§4).
+(3) Empirical insights into human-LLM gaps.
+Our analyses reveal that scholarly writing is highly non-linear and multi-intentional: more than half of all writing sessions involve three or more intertwined intentions.
+Writers dedicate most time to producing and revising text for clarity and coherence, while cognitively heavy tasks, such as idea organization or visual structuring, occur in fewer but longer, more focused sessions.
+Evaluations of current LLMs (GPT-5, Qwen) show they can imitate surface-level edits but fail to predict next intentions or sustain cognitively complex revisions, underscoring the limits of current models as cognitively aligned writing partners (§5 and §6).
 2 Related Work
-2.1 Cognitive Theory of Writing Process
-Research on human writing shifted from analyzing final written products to examining cognitive processes across various writing phases Diederich (1974); Krapels (1990); MacArthur and Graham (2016).
-Building on this process-oriented approach, Flower and Hayes (1981)’s cognitive writing theory outlines three key sub-processes: (1) planning, (2) translating, and (3) reviewing.
-These interconnected and non-linear stages of human writing inform our work, where we expand this model to a more granular taxonomy of cognitive writing patterns in scholarly communication.
+Cognitive Theory of Writing Process
+Research on writing has shifted from analyzing final texts to examining cognitive processes across writing phases (Diederich, 1974; Krapels, 1990; MacArthur and Graham, 2016).
+Building on this process-oriented approach, Flower and Hayes (1981) outline three core sub-processes: planning, translating, and reviewing.
+These dynamic, non-linear stages inform our work, where we extend this model into a finer-grained taxonomy of cognitive writing patterns in scholarly communication.
 Koo et al.
-(2023) introduced a taxonomy of scholarly writing process based on keystroke data from short research plans written in 30-minute sessions.
-Expanding on this, we perform a larger-scale keystroke collection over months, whose final products culminated in research publications.
-Expert-reviewed and grounded in prior literature, our taxonomy encompasses end-to-end cognitive trajectories of scholarly writing.
-2.2 Keystroke Loggers for Scholarly Writing
-Keystroke logging tools (e.g., Inputlog) allow researchers to observe digital writing without disrupting the writing process Chan (2017); Johansson et al.
-(2010); Leijten and Van Waes (2013); Lindgren and Sullivan (2019).
-However, they are often restricted to a closed ecosystem, such as MS Word, and thus less accessible for scientific communities who use LaTeX.
-Moreover, they are not well-suited for collecting data from extended writing sessions.
-Current studies rarely examined scholarly communication that uses LaTeX, particularly for scientific writing.
-To address this gap, we developed systems that securely collect real-time keystrokes over months from Overleaf, a widely-used online LaTeX editor, while ensuring privacy.
-The system workflow enables span-level annotation of writing intentions, offering a natural, uninterrupted environment for studying cognitive writing activities over long periods and across multiple sections.
-2.3 Datasets for Scholarly Writing
-Previous work primarily focused on constructing datasets to analyze scholarly writing processes, which vary in content and scale.
-Publicly available datasets tend to track linguistic style changes or grammatical edits during revision Du et al.
-(2022b); Jiang et al.
-(2022); Ito et al.
-(2019); Mita et al.
-(2022), while others capture edits based on feedback and peer review D’Arcy et al.
-(2024); Jourdan et al.
-(2024); Kuznetsov et al.
-(2022) or citation generation Kobayashi et al.
-(2022); Narimatsu et al.
-(2021).
-Also, most focus on specific sections of papers (e.g., abstracts, introductions) Du et al.
-(2022b); Mita et al.
-(2022).
-In contrast, our dataset covers all cognitive phases of writing, from inception to final manuscript, over an extended period.
-Furthermore, those existing corpora often compare multiple versions of final manuscripts from preprint databases Jiang et al.
-(2022); Du et al.
-(2022b); D’Arcy et al.
-(2024); Jourdan et al.
-(2024), missing how those manuscripts evolved Jourdan et al.
-(2023).
-To address this, we build a keystroke-based corpus that captures the progression of publications in real-time (See Figure 2).
-3 Dataset Construction
-To collect authentic writing data from writers’ natural thoughts, we developed a Chrome extension that records keystroke data in the background.
-After the study, an in-depth annotation process is conducted to interpret the intent behind each keystroke.
-We describe the detailed process of our data collection, annotation system design, and taxonomy creation in the following sections.
+(2023) proposed a taxonomy of scholarly writing based on keystroke data from short, 30-minute research-plan sessions.
+Expanding on this, we collect larger-scale, longitudinal keystroke data spanning months and culminating in published research.
+Expert-reviewed and grounded in prior theory, our taxonomy captures end-to-end cognitive trajectories of scholarly writing.
+Keystroke Loggers for Scholarly Writing
+Keystroke logging tools (e.g., Inputlog) enable unobtrusive observation of digital writing (Chan, 2017; Johansson et al., 2010; Leijten and Van Waes, 2013; Lindgren and Sullivan, 2019).
+Yet, most operate in closed ecosystems like MS Word and are ill-suited for LaTeX-based academic writing or extended sessions.
+To address this, we built systems that securely record real-time keystrokes from Overleaf—an online LaTeX editor—over months while preserving privacy.
+This workflow supports span-level annotation of writing intentions, providing a natural setting to study long-term cognitive writing activities across document sections.
+Datasets for Scholarly Writing
+Publicly available datasets from previous work tend to track linguistic style changes or grammatical edits during revision (Du et al., 2022b; Jiang et al., 2022; Ito et al., 2019; Mita et al., 2022), while others capture edits based on feedback and peer review (D’Arcy et al., 2024; Jourdan et al., 2024; Kuznetsov et al., 2022) or citation generation (Kobayashi et al., 2022; Narimatsu et al., 2021).
+Also, most focus on specific sections of papers (e.g., abstracts, introductions) (Du et al., 2022b; Mita et al., 2022).
+In contrast, our dataset covers all cognitive phases of writing over an extended period and all sections.
+Furthermore, those existing corpora often compare multiple versions of final manuscripts from preprint databases (Jiang et al., 2022; Du et al., 2022b; D’Arcy et al., 2024; Jourdan et al., 2024), missing how those manuscripts evolved (Jourdan et al., 2023).
+To address this, we build a keystroke-based corpus that captures real-time progression of publications.
+3 Data Collection and Intent Annotation
+To capture scholarly writing as it naturally unfolds, we developed a system that records the end-to-end writing process, from individual keystrokes to their cognitive interpretation, without interrupting authors’ workflow.
+This section introduces the design of our Chrome extension for Overleaf-based data collection, the participant recruitment protocol, and the annotation pipeline that underpins our taxonomy of cognitive writing intentions.
+3.1 Keystroke Collection in Overleaf
+We built a custom Chrome extension (Appendix Figure 8) that unobtrusively logs real-time keystroke trajectories from Overleaf.
+The extension operates in the background once participants consent and authenticate via unique credentials.
+Every time a key-up event occurs, the system captures the visible text within the user’s Overleaf editor, compares it with the previous version using Google’s diff_match_patch algorithm, and stores the resulting text differences alongside metadata such as timestamp, file name, and author ID.
+To protect privacy, only project IDs pre-approved through participant consent were collected; any non-consented Overleaf projects were automatically filtered out.
+All data were stored on a secure institutional server and anonymized before analysis.
+Appendix A.2 provides further implementation details.
 3.2 Participant Recruitment
-We recruited 10 graduate students222All participants attend a R1 university in US and are proficient in English.
-Appendix A.1 details the recruitment process.
-from the computer science department who were actively preparing academic manuscripts in Overleaf LaTeX editor for submission to peer-reviewed conferences in AI and NLP.
-The data is collected from November 2023 to February 2024, totaling up to 4 months.
-Our study has been approved by the IRB institution of the authors.
-Please refer to Appendix A.1 for more detailed descriptions.
-3.2 Chrome Extension for Keystroke Collection in Overleaf
-We designed and implemented a Chrome extension (Appendix Figure 7), which enables the real-time collection of keystroke trajectories in the Overleaf platform.
-Participants can create their account credentials, and after logging into the system, the extension monitors the user’s keystrokes in the background silently without disrupting writing process.
-See Appendix A.2 for more description of the extension workflow.
+We recruited 10 graduate students in computer science at an R1 university in the United States, each actively writing manuscripts for peer-reviewe AI and NLP conferences using Overleaf.
+The collection period spanned four months (November 2023–February 2024), capturing writing sessions over the full lifecycle of manuscript preparation.
+All participants were proficient in English and provided informed consent under IRB-approved protocols (Appendix A.1).
+Each participant installed the Chrome extension on their personal device and continued their usual writing activities, allowing us to gather naturalistic data without external disruption.
+3.3 Writing Intention Annotation
+To uncover the cognitive mechanisms underlying scholarly writing, we annotate every collected keystroke with its corresponding writing intention.
+This annotation process transforms raw writing traces into interpretable cognitive data, forming the foundation of the ScholaWrite taxonomy and subsequent analyses.
+Each keystroke record contains metadata, such as file name, type of action, text differences between two states, and line number in the Overleaf editor, which allows precise reconstruction of how writers iteratively develop and refine their manuscripts.
+Please see Appendix B for further annotation details.
+Annotation Interface
+We developed an interactive annotation interface that visualizes writing activity over time and across authors and files (Figure 9 in Appendix).
+Annotators can navigate a timeline of keystrokes, examine spans of related edits (e.g., drafting a sentence or clarifying an argument), and assign one or more intention labels from a predefined taxonomy via a dropdown menu.
+This design allows annotators to interpret not only what changes were made, but why they occurred, linking textual behavior to underlying cognitive intent.
 | 1st | Intention | Definition | An example action | Prop. |
-Planning | Idea Generation | Formulate and record initial thoughts and concepts. | writing keywords or notes (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
+| Planning | Idea Generation | Formulate and record initial thoughts and concepts.
+| writing down keywords of a paragraph beforehand (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
 | Idea Organization | Select the most useful materials and demarcate those generated ideas in a visually formatted way.
@@
 […]
-[…]
-Coherence | Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole.
+| Revision | Fluency | Fix grammatical or syntactic errors in the text or LaTeX commands. | “We desiginingned several experiment setups for the LLM evaluations as described..” | 1.4% |
+| Coherence | Logically link (1) multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent.
 | “Each comment was annotated by three different annotators , which and we achieved high inter-annotator agreement.” | 3.3% | |
@@
 | Structural | Improve the flow of information by modifying the location of texts and objects.
-| “We calculate Pearson’s correlation for human alignment to compare the alignment between lexicon-based preferences and humans’ original preferences.
-First, we calculated the score from each human participant to compare the alignment between lexicon-based preferences and humans’ original preferences.” | 3.7% | |
+| “..human alignment to compare the alignment between lexicon-based preferences and humans’ original preferences.
+First, we calculated the score from each human participant to compare the alignment between..” | 3.7% | |
 | Linguistic Style | Modify texts with the writer’s writing preferences regarding styles and word choices, etc. | “We believe posit that …” | 1.6% | |
@@
 […]
-3.3 Writing Intention Annotation
-To decode the nuanced intentions behind each writing action, we annotate each keystroke collected from the Chrome extension, using a specialized interface.
-This process is crucial for developing a detailed taxonomy of the end-to-end scholarly writing process, which serves as the basis for annotating the collected keystrokes.
-To perform annotations, we processed those raw keystroke entries by file name, type of writing actions, an array of differences between two subsequent texts, and line number in an Overleaf editor.
-Please see Appendix B for further annotation details.
-Annotation Interface
-We developed a novel interface for keystroke annotation, enabling visualization by time, LaTeX file, and author (Appendix Figure 8).
-Annotators can navigate the keystroke timeline, identify a span of related edits (e.g., outlining or grammar correction), and assign a pre-defined intention label from the dropdown menu.
 Intention Taxonomy Construction
-Our taxonomy (Table 1) builds on prior literature Flower and Hayes (1981); Du et al.
-(2022b); Koo et al.
-(2023).
-During annotations, we have identified 15 types of scholarly writing intentions from the five projects and further grouped them into three high-level categories according to Flower and Hayes (1981) - Planning, Implementation, and Revision.
+Building on cognitive theories of writing (Flower and Hayes, 1981) and recent empirical studies on scholarly revision (Du et al., 2022b; Koo et al., 2023), we developed a taxonomy of 15 distinct writing intentions grouped into three overarching categories (Table 1).
 Following Pustejovsky et al.
-(2017), we performed an iterative open coding for the collected keystrokes from the two Overleaf projects.
-Using the annotation interface, each annotator reviewed the first few writing keystrokes and identified the high-level processes.
-Next, the annotators (1) determine the span of keystrokes that form a meaningful unit together (e.g., a clause or sentence) within the high-level process and (2) decide the underlying intention of edits.
-The annotators coded these spans independently, resolving conflicts through multiple in-depth discussions.
-Finally, a cognitive linguist participated in all discussions, refining the definition and scope of each writing intention.
-To ensure taxonomy validity, we followed two standard practices Nickerson et al.
-(2013); Kundisch et al.
-(2021): (1) mutual exclusivity and (2) collective comprehensiveness.
-Ambiguous keystroke spans were flagged and refined through iterative discussions to sharpen category boundaries.
-After finalizing the codebook and taxonomy, we updated intention annotations333The updated list appears in annotation interface.
-for the two selected projects and applied the taxonomy to the rest.
-The final inter-annotator agreement was (weighted F1 score in a multi-label setting for a 1K-keystroke sample444Unlike Kappa, which assumes categorical agreement and struggles with multi-label classification, weighted F-1 scores better handle class imbalances and partial agreements.).
-See Appendix B for full annotation details.
-4 ScholaWrite: A Dataset of Cognitive Process of Scholarly Writing
-Summary
-Our post-processed dataset, ScholaWrite, contains end-to-end writing trajectories with annotated intentions, for five Overleaf projects whose final product turned into arXiv preprints.
-The dataset consists of 61,504 arrays of keystroke changes with fine-grained annotations of writing intentions.
-Appendix Table 6 shows the overall statistics of the dataset.
-Writing Intention Distributions
-The writing intention distributions is illustrated in Figure 3(c).
-During the planning stage, idea generation is the most frequent intention.
-Text production dominates the implementation stage, and clarity becomes the primary focus during the revision stage.
-Since we defined the boundary between implementation and revision as the point where meaning changes occur, the high frequency of text production suggests that authors often continue revising their texts by introducing new topics into the documents.
-Appendix Table 8 shows the label distribution per project.
-Label | Subsequent label | Probability | |
-Idea Generation | Text Production | 0.52 | |
-Idea Organization | Idea Generation | 0.34 | |
-Citation Integration | Text Production | 0.37 | |
-Cross-reference | Text Production | 0.36 | |
-Clarity | Text Production | 0.35 | |
-Coherence | Text Production | 0.34 | |
-Scientific Accuracy | Text Production | 0.34 |
-Flows Between Writing Intentions
-We analyze the flow of writing intentions most likely to follow each other throughout the writing process, as presented in Table 2.
-Many intentions proceed to text production, where the idea generation has the highest probability of feeding into text production.
-Figure 3(a) also shows a similar pattern of flows between text productions and other intentions.
-Also, in Table 2 the reflexive relationship between text production and clarity suggests a constant loop of producing new texts and refining them, ensuring the precision and reasonability of their claims.
-Please see Appendix Figure 13 for the intention flows of each of the five projects.
-Time-series Distributions of Writing Intentions
-We calculated the average Wasserstein distance between each intention distribution and uniform distribution to assess how evenly intentions occur throughout the writing process.
-As shown in Figure 3(d), text production had the smallest distance, while scientific accuracy had the largest.
-This suggests that text production is spread throughout the writing, whereas scientific accuracy is concentrated in specific phases.
-A similar pattern appears in Figure 3(e) , where text production is evenly distributed over time, and scientific accuracy peaks in the middle-to-late of the writing process.
-Figure 3(b) demonstrates a similar pattern that text production of the implementation stage spreads evenly over time.
-Types of clarity and structural revisions tend to occur after the medium steps of writing.
-Please see Appendix Figures 9 to 12 for each of the five projects.
-5 Applications for Writing Assistance
-We envision ScholaWrite as a valuable resource for training language models and improving future writing assistants for scholarly writing.
-To evaluate its usability, we conducted experiments training LLMs to mimic the complex, non-linear writing processes of human scholars.
-Specifically, we aimed to showcase the capabilities of LLMs trained on ScholaWrite in two scenarios:
-(1) Predicting an author’s next writing intention (§5.1): The task is crucial for writing assistants to accurately assess the writer’s current status in context and predict the correct writing intention.
-This enables them to offer cognitively-appropriate writing suggestions that align with the writer’s needs.
-(2) Iteratively generating scholarly writing actions from scratch (§5.2) (called Iterative Self-Writing), mirroring the human writing process: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
-5.1 Predicting Next Writing Intention
-Setup & Metrics
-This task (represented by the inner box “Prediction” in Figure 4) takes the “before-text” from a keystroke pair (i.e., Listing LABEL:table:single-entry) and context prompt, and predicts the writing intention to apply for the subsequent actions.
-We use BERT Devlin et al.
-(2019), RoBERTa Liu et al.
-(2019), and Llama3.1-8B-Instruct Dubey et al.
-(2024) as baselines, fine-tuning each on the ScholaWrite training set.
-For comparison, we also run GPT-4o OpenAI (2024) on the test set.
-To evaluate model performance, we used a weighted F-1 score555Weighted F-1 was chosen to address skewed label distribution (as shown in Figure 3(c) and Table 1)..
-See Appendix E.2 for details.
+(2017), two annotators conducted iterative open coding on initial subsets of keystrokes to inductively identify recurring cognitive patterns.
+Each span of edits was treated as a meaningful writing unit, such as composing a sentence or improving phrasing, and assigned an appropriate intention label.
+Disagreements were resolved through multiple rounds of discussion with a cognitive linguist, who refined label definitions and boundaries to ensure coherence.
+To validate the taxonomy, we applied two principles from taxonomy design (Nickerson et al., 2013; Kundisch et al., 2021): mutual exclusivity (each edit corresponds to one primary intention) and collective comprehensiveness (taxonomy covers all observed behaviors).
+After iterative refinement, we achieved a weighted F1 inter-annotator agreement of on a 1K-keystroke subset, confirming strong reliability.
+See details in Appendix B.
+Post-processing
+After annotation, we post-processed the data to improve usability and ensure privacy.
+All personal identifiers were removed, and keystrokes labeled as non-informative artifacts (e.g., scrolling, idle typing) were filtered out.
+Remaining entries were segmented into discrete intention spans suitable for analysis and model training.
+(See details in Appendix A.3).
+4 ScholaWrite
+ScholaWrite represents the first large-scale publicly available dataset capturing the cognitive dynamics of end-to-end scholarly writing.
+It links every keystroke to the author’s underlying writing intention, allowing direct observation of how ideas evolve into finalized text.
+Each annotated entry follows a standardized schema connecting text changes, contextual metadata, and intention labels, as shown below, illustrating a Fluency edit, in which the writer corrects “expct” to “aspect.”
+4.1 Overview
+The final ScholaWrite dataset contains 61,504 keystroke-based writing actions collected across five Overleaf projects, each of which culminated in an arXiv preprint authored by graduate researchers.
+Every action is annotated with a cognitive writing intention drawn from the 15-category taxonomy described in Section 3.3.
+Table 2 summarizes key statistics across projects.
+In total, the dataset captures over 62K text edits, encompassing more than 118K words added or deleted across multi-month writing trajectories.
+Each project contains thousands of temporally ordered keystrokes, providing a detailed chronicle of the writing process at the sentence, paragraph, and document levels.
+4.2 Analytical Units: Capturing Writing Flow
+Because the dataset records precise timestamps in milliseconds, we can reconstruct the temporal structure of scholarly writing.
+To analyze this structure, we define two complementary units: intention sessions and sessions.
+Intention session.
+An intention session is a continuous period during which the writer focuses on a single cognitive goal, such as producing text, improving clarity, or inserting a figure, before switching to another intention.
+These sessions typically last from a few seconds to several minutes, capturing a concentrated burst of focused cognitive activity.
+We apply two constraints for meaningful segmentation: (1) inactivity longer than 10 minutes marks the end of a session; (2) sessions shorter than 30 seconds are discarded to exclude trivial edits.
+This definition allows us to quantify how long writers sustain specific cognitive modes before transitioning to new tasks.
+| Project | 1 | 2 | 3 | 4 | 5 |
+| # Authors | 1 (3) | 1 (4) | 1 (3) | 1 (4) | 9 (18) |
+| # Keystrokes | 14,217 | 5,059 | 6,641 | 8,348 | 27,239 |
+| # Words added | 17,387 | 23,835 | 7,779 | 12,448 | 57,511 |
+| # Words deleted | 11,739 | 15,158 | 2,308 | 7,621 | 25,853 |
+Session.
+In contrast, a session aggregates all writing actions within a continuous working period, regardless of how many intentions are involved.
+It represents the natural rhythm of scholarly writing, how authors weave together planning, drafting, and revising while working on a particular section or idea.
+Sessions end after 10 minutes of inactivity and must last at least one minute.
+5 Analyses
+ScholaWrite offers an unprecedented window into the dynamic, cognitive process of scholarly writing.
+Our analysis pursues two complementary goals: (1) to uncover how scholarly writing intentions evolve, interact, and shift across time, and (2) to inform the design of cognitively aligned writing assistants capable of supporting real-world research workflows.
+We focus on three research questions:
+5.1 Which Writing Intentions Dominate?
+We first examine which of the fifteen annotated writing intentions account for the greatest share of time and cognitive effort.
+| Intention | % Time | Intention | % Time | ||
+| Idea Generation | 4.8 | Fluency | 3.3 | ||
+| Idea Organization | 0.6 | Coherence | 6.9 | ||
+| Section Planning | 2.6 | Clarity | 13.4 | ||
+| Text Production | 40.3 | Structural | 4.0 | ||
+| Object Insertion | 8.4 | Linguistic Style | 2.6 | ||
+| Citation Integration | 2.4 | Scientific Accuracy | 1.2 | ||
+| Cross-reference | 2.4 | Visual Formatting | 6.1 | ||
+| Macro Insertion | 0.8 |
+Time Distribution across Intentions
+Table 3 summarizes the average share of time spent per intention across five projects.
+Text Production dominates with 40.3% of total time, followed by Revision intentions – mainly Clarity, Coherence, and Visual Formatting – at 26.4%.
+In contrast, Planning activities (e.g., Idea Generation, Organization) take less than 5%.
+These results suggest that scholars devote most of their time to articulating and refining text rather than explicit planning.
+The considerable effort spent on revision and formatting underscores that much of scholarly writing involves clarifying and visually shaping ideas, an aspect often overlooked in computational models of writing.
+Session Duration and Cognitive Effort
+We analyze the average duration of each intention session – a continuous period focused on one cognitive goal.
+Table 4 shows that Idea Organization sessions last the longest (3.7 min), followed by Object Insertion (2.9 min) and Visual Formatting (2.2 min), all showing high variability due to complex tasks like reorganizing structure or managing visuals.
+In contrast, Text Production, though dominant in total time, consists of short, frequent bursts (mean 2.2 min).
+This contrast reveals two cognitive modes: (1) short, frequent drafting episodes, and (2) long, focused planning or visual composition that require sustained attention.
+| Intention | Mean | SD | Max | Count |
+| Idea Generation | 1.87 | 1.94 | 11.40 | 76 |
+| Idea Organization | 3.71 | 4.58 | 13.79 | 7 |
+| Section Planning | 1.37 | 1.57 | 9.36 | 38 |
+| Text Production | 2.20 | 2.53 | 28.14 | 724 |
+| Object Insertion | 2.88 | 3.05 | 21.61 | 108 |
+| Citation Integration | 1.70 | 1.53 | 7.53 | 34 |
+| Cross-reference | 1.08 | 0.77 | 3.98 | 38 |
+| Macro Insertion | 1.43 | 1.05 | 3.71 | 7 |
+| Fluency | 0.66 | 0.18 | 1.06 | 10 |
+| Coherence | 1.46 | 1.71 | 9.62 | 55 |
+| Clarity | 1.64 | 1.61 | 9.53 | 243 |
+| Structural | 1.50 | 1.43 | 8.42 | 53 |
+| Linguistic Style | 2.05 | 2.55 | 10.81 | 24 |
+| Scientific Accuracy | 1.97 | 1.60 | 6.44 | 20 |
+| Visual Formatting | 2.22 | 2.66 | 14.25 | 87 |
+5.2 How Are Writing Intentions Intertwined?
+Scholarly writing rarely unfolds linearly.
+Instead, it involves recursive alternations between complementary intentions, writers generate ideas, structure them, produce text, and immediately refine it.
+As seen in Figure 1 (and other projects in Appendix Figure 14), the intertwined processes of writing reveals dense bidirectional flows across intentions.
+Transition Patterns
+The transition probability matrix (Figure 2) shows that more than half of all Planning actions (e.g., Idea Generation) are immediately followed by Text Production.
+Conversely, Idea Organization often leads back to Idea Generation, suggesting that structuring ideas can prompt new content creation.
+Within Implementation, most transitions also return to Text Production (>30%), except for Macro Insertion, which frequently precedes Idea Generation.
+In Revision, strong recursive loops emerge, especially among Clarity, Fluency, and Coherence, highlighting how writers iteratively refine language and argumentation.
+Overall, the results indicate that planning, implementation, and revision are not distinct stages but interdependent cycles continually revisited throughout the writing process.
+Multitasking and Alternation Patterns
+Writing sessions are highly multitasking in nature.
+As shown in Figure 3, 57% of all sessions involve three or more distinct intentions, reflecting frequent switching between cognitive modes within a single sitting.
+This dynamic interplay reveals how writers manage multiple goals in parallel.
+Frequent 6-gram intention sequences (Appendix Table 9) confirm this recursive pattern.
+The most common alternating pairs include: (Text Production, Clarity), (Text Production, Object Insertion), and (Idea Generation, Idea Organization).
+These short feedback loops illustrate the micro-structure of scientific writing: writers repeatedly produce, adjust, and reorganize – blending idea formation with textual execution.
+| Intention | Early | Late | Trend |
+| Idea Gen. | 1.6 (23) | 2.5 (11) | longer, fewer |
+| Idea Org. | 2.6 (2) | — | early-only |
+| Sect. Plan. | 1.6 (13) | 1.0 (17) | more, shorter |
+| Text Prod. | 2.1 (221) | 2.0 (227) | stable |
+| Obj. Insert. | 2.4 (28) | 3.4 (42) | up both |
+| Cite Int. | 1.4 (16) | 1.8 (6) | fewer later |
+| Cross-ref. | 0.7 (8) | 1.2 (18) | more later |
+| Clarity | 1.5 (43) | 1.3 (99) | more, quicker |
+| Structural | 2.4 (8) | 1.4 (29) | more later |
+| Ling. Style | 3.5 (3) | 1.7 (12) | concise edits |
+| Vis. Form. | 3.6 (16) | 1.7 (43) | quick fixes |
+5.3 Are writing patterns stable throughout the process, or do they change across phases?
+Scholarly writing is not static, it evolves as authors move from idea formation to refinement.
+Beyond the aggregate trends observed earlier, we examine how writing intentions shift across time by comparing the early phase (first third of project duration) with the late phase (final third).
+Which intents dominate at different stages?
+Figure 4(a) shows how writing time is allocated across phases.
+Planning intentions are more frequent early (9.1% vs.
+5.4%), reflecting front-loaded structuring.
+Within Implementation, Citation Integration is higher early, while Cross-referencing increases late as authors add visuals and refine internal links.
+Revision grows sharply in the late phase (46.7% vs.
+29.6%), consistent with end-stage polishing for coherence and presentation.
+Table 5 summarizes session-level differences.
+In Planning, Idea Generation and Organization sessions decrease in count but lengthen, suggesting deeper cognitive engagement later.
+Section Planning rises as writers reorganize material for submission.
+For Revision session counts increase substantially, though durations shorten, indicating rapid, localized edits characteristic of final revisions.
+How are intentions intertwined across stages?
+Figure 15 in Appendix compares transition probabilities across phases.
+Early writing features strong flows into Implementation, whereas late writing shifts toward Revision, marking the transition from building content to refining it.
+Some intention paths also evolve — for example, transitions from Macro Insertion lead to Idea Generation early but to Section Planning late, as layout decisions become more prominent.
+Writers also juggle more intentions later in the process.
+As shown in Figure 4(b), 19.4% of late sessions include over five distinct intentions (vs.
+14.8% early), reflecting increased multitasking and integration of planning, drafting, and revising.
 | BERT | RoBERTa | Llama-8B | GPT-4o | |
@@
 | + SW | 0.64 | 0.64 | 0.13 | - |
-Iter. | Llama-8B-SW | Llama-8B-Zero | GPT-4o |
-[…]
-[…]
-[…]
-[…]
-Results
-Table 3 presents the weighted F1 scores for predicting writing intentions across baselines and fine-tuned models.
-Regardless of the intricate nature of the task itself666Each model predicts the next intention using only the ”before” text, while human annotators consider multiple ”before and after” edits.
-Moreover, the chosen next intention is not necessarily the only correct one., all models finetuned on ScholaWrite show an improved performance compared to their baselines.
-BERT and RoBERTa achieved the most improvement, while LLama-8B-Instruct showed a modest improvement after fine-tuning.
-As detailed in Appendix E, more training epochs and encoder-decoder architectures of BERT variants are assumed to be the reason for significant improvement compared to LLMs.
-This aligns with findings from Grasso and Locci (2024); Yu et al.
-(2023), which shows that RoBERTa and BERT can often match or even outperform LLMs for the text classification tasks.
-Those results demonstrate the effectiveness of our ScholaWrite dataset to align language models with writers’ intentions.
-5.2 Iterative Self-Writing
-Setups
-During iterative self-writing (Figure 4), a fine-tuned model processes LaTeX-formatted seed document (as “before-text”) with a context prompt to predict the next intention, then revises the text (“after-text”) accordingly given prompt.
-The revised document then serves as the new seed for the next iteration.
-This process repeats until a set iteration limit (e.g., 100) is reached.
-All models use the same train (80%)-test (20%) split across experiments.
-See Appendix E and Figure 14 for training details.
-We fine-tune Llama3.1-8B-Instruct (Llama-8B-SW) and compare it to vanilla Llama-8B-Instruct (Llama-8B-Zero) and GPT-4o.
-Also, seed documents were derived from LaTeX-formatted abstracts of four award-winning NLP papers on diverse topics Zeng et al.
-(2024); Lu et al.
-(2024b); Du et al.
-(2022a); Etxaniz et al.
-(2024) (Appendix LABEL:table:seed-entry-johnny-LABEL:table:seed-entry-latxa).
-Metrics
-We evaluated lexical diversity (unique-to-total token ratio), topic consistency (cosine similarity between seed and final output), and intention coverage (unique writing intentions used over 100 iterations).
-For human evaluation, three native English speakers with LaTeX expertise assessed outputs from Llama-8B-Zero and Llama-8B-SW on accuracy (alignment with predicted intention), alignment (similarity to human writing), fluency(grammatical correctness), coherence(logical structure), and relevance (connection to the seed document) - Refer to Appendix LABEL:sec:appendix:human-eval.
-Accuracy was judged per iteration, while other metrics used pairwise comparisons.
-Inter-annotator agreement (IAA)777The IAA scores are 0.84 (SW) and 0.76 (Zero) for accuracy, all 100% for alignment, fluency, coherence, and 49.8% (SW) and 100% (Zero) for relevance.
-was measured using Krippendorff’s alpha for accuracy and percentage agreement for others.
-Metrics | Model | Seed 1 | Seed 2 | Seed 3 | Seed 4 |
-Accuracy | SW | 21.0 | 10.3 | 18.3 | 15.3 |
-Zero | 35.7 | 29.7 | 45.3 | 43.3 | |
-Alignment | SW | 0 | 0 | 0 | 0 |
-Zero | 3 | 3 | 3 | 3 | |
-Fluency | SW | 0 | 0 | 0 | 0 |
-Zero | 3 | 3 | 3 | 3 | |
-Coherence | SW | 0 | 0 | 0 | 0 |
-Zero | 3 | 3 | 3 | 3 | |
-Relevance | SW | 1 | 3 | 2 | 1 |
-Zero | 3 | 3 | 3 | 3 |
-Results
-Figure 5 shows that Llama-8B-SW consistently produced the most lexically diverse words, generated the most semantically aligned topics (Seeds 1 & 2), and covered the most writing intentions (except Seed 3).
-These results underscore the value of ScholaWrite in improving scholarly writing quality generated by language models.
-However, our human evaluation (Table 5) revealed that Llama-8B-SW generated less human-like writing, in terms of fluency and logical claims.
-It also struggled with generating texts aligned with the predicted intentions.
-See Appendix Tables LABEL:table:human-eval-seed1 to LABEL:table:human-eval-seed4 for more details.
-Despite the weaknesses, Llama-8B-SW still produced more relevant content (Seed 2), which aligns with topic consistency trends in Figure 5, highlighting the usefulness of ScholaWrite dataset in certain contexts.
-Moreover, Llama-8B-SW exhibited the most human-like writing activity patterns over time (Figure LABEL:fig:writing-step-intention-all-model), which frequently switches between implementation and revision and covers all three high-level processes.
-Llama-8B-Zero and GPT-4o tend to remain in a single high-level stage throughout all 100 iterations of self-writing.
-Compared to Appendix Figure 12, which depicts frequent transitions across all three stages in an early draft (e.g., the first 100 steps), Llama-8B-SW most closely replicates human writing behaviors in iterative writing tasks.
-These findings reinforce the potential of ScholaWrite in helping LLMs emulate human scholarly writing processes.
+6 Can LLMs Support Human Writing?
+Scientific writing is deeply cognitive, non-linear, and phase-dependent.
+Writers continuously shift between intents, and multitasking across intertwined intentions.
+For LLMs to genuinely assist this process, they must (i) infer what the writer is doing or about to do, and (ii) provide support that aligns with the writer’s actual cognitive intent.
+We evaluate these capabilities through two tests using ScholaWrite:
+6.1 Predicting Next Writing Intention
+Because writing involves frequent task-switching, an effective assistant should anticipate the writer’s next move.
+We test whether models can predict the upcoming writing intention based on the current text state.
+Setup.
+We use the “before-text” from each keystroke pair and prompt the model to predict the next writing intention.
+Models evaluated include BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), Llama3.1-8B-Instruct (Dubey et al., 2024), and GPT-4o, with fine-tuned versions train on our dataset.
+Performance is measured using weighted F1 to account for label imbalance.
+See details in Appendix G.2.
+Findings.
+Base models perform poorly (F1 0.12), indicating that current LLMs cannot infer writing intent from text alone.
+Fine-tuning on ScholaWrite improves performance dramatically (to 0.64 for BERT/RoBERTa), showing that cognitively annotated writing data helps models anticipate human writing behavior, a key step toward adaptive, context-aware assistance.
+Training details and explanations for differences across models are reported in Appendix F.1.
+6.2 Output Alignment across Writing Sessions
+To assist writers effectively, a model must not only infer their intentions but also generate text aligned with those goals.
+We evaluate LLMs under two settings: (1) single-intention sessions, focused on one cognitive goal, and (2) multi-intention sessions, where multiple intentions intertwine.
+Setup.
+For each session, models receive the “before-text” and either a single intention label or the sequence of intentions observed in that session, then generate the corresponding “after-text.” We use GPT-5 (OpenAI, 2025) as the base model and evaluate alignment with human writing using BERTScore-F1 (Zhang et al., 2019).
+Results for other models (e.g., Qwen (Yang et al., 2024)) and metrics (e.g., Levenshtein distance) appear in the Appendix F.2 and show consistent patterns.
+Findings.
+Model alignment decreases as sessions grow longer or cognitively more complex (Figure 5).
+As shown in Figure 6, GPT-5 (and Qwen) perform best on Revision intentions (e.g., Clarity, Fluency) but struggle with Planning and Implementation, which demand deeper reasoning and structural organization.
+Alignment further drops when multiple intentions co-occur within a session, indicating models’ limited ability to integrate overlapping cognitive goals.
+Overall, current LLMs excel at surface refinement but remain weak at sustaining the intertwined reasoning processes that drive human scholarly writing.
 7 Conclusion and Future Work
-We present ScholaWrite, a first-of-its-kind dataset capturing the end-to-end cognitive process of scholarly writing, comprising nearly 62K LaTeX keystrokes collected via our custom-built Chrome extension.
-This dataset, sourced from ten graduate students with varying levels of scientific writing expertise, is further enriched with expert annotations based on a novel taxonomy of cognitive writing intentions inspired by Flower and Hayes (1981); Koo et al.
-(2023).
-Through several experiments, ScholaWrite shows its value for advancing the cognitive capabilities of LLMs and developing cognitively-aligned writing assistants, enabling them to mimic the complex, non-linear, and intention-driven nature of human writing.
-Future work includes expanding the dataset to diverse academic fields, authors, and collaborative projects, thus enabling models to generalize beyond fact recall to emulate human decision-making and reasoning in more realistic academic environments.
-Additionally, integrating advanced memory architectures and lifelong learning techniques could further enhance LLMs’ ability to adapt dynamically to evolving writing intentions and produce coherent, high-quality scholarly outputs.
+This work introduces ScholaWrite, the first dataset to capture the end-to-end cognitive process of scholarly writing.
+Unlike prior work that centers only on final manuscripts, ScholaWrite reveals how writing unfolds through iterative, multi-intent, and phase-dependent behaviors.
+Our analyses uncover three key insights.
+First, scholarly writing is highly non-linear: authors continually alternate between planning, implementing, and revising, rather than progressing in discrete stages.
+Second, writing intentions evolve across phases, with early stages emphasizing idea development and late stages dominated by refinement and multitasking.
+Third, current LLMs, though proficient at surface-level revision, struggle to anticipate or integrate complex, cognitively demanding intentions, revealing a critical gap between human writing cognition and model capabilities.
+ScholaWrite contributes: (1) a fine-grained lens on the cognitive dynamics of scholarly writing; (2) a benchmark for evaluating LLMs’ alignment with human writing intentions; and (3) a training resource for developing models that adapt to evolving, multi-intent writing behavior.
+Our Chrome extension enables unobtrusive, in-situ data capture in Overleaf, demonstrating a scalable and ethical method for collecting realistic writing data.
+Future work will extend beyond keystroke logs to include pre-writing ideation, collaboration, and multimodal composition (e.g., figure and data integration).
+By tracing the full scholarly writing lifecycle, we aim to deepen our understanding of human cognitive patterns and to design AI writing assistants that not only complete text, but truly complement how humans think and write.
 Limitations and Ethical Considerations
@@
 While the number of articles is limited, the granularity and volume of the data provide a rich resource for understanding writing behaviors.
-Prior research has shown that detailed keystroke logs, even from small datasets, can effectively model writing processes Leijten and Van Waes (2013); Guo et al.
-(2018); Vandermeulen et al.
-(2023).
+Prior research has shown that detailed keystroke logs, even from small datasets, can effectively model writing processes (Leijten and Van Waes, 2013; Guo et al., 2018; Vandermeulen et al., 2023).
 Unlike studies focused on final outputs, our dataset enables a process-oriented analysis, emphasizing the cognitive and behavioral patterns underlying scholarly writing.
@@
 Expanding to multilingual settings could reveal unique cognitive and linguistic insights into writing across languages.
-Fourth, due to computational and cost constraints, we evaluated the usability of the ScholaWrite dataset with a limited number of LLMs and hyperparameter configurations.
-As shown in Table 3, the Llama-8B-Instruct model demonstrated only marginal improvements after fine-tuning on our dataset.
-This underscores the need for future research to explore advanced techniques, such as fine-grained prompt engineering, to better align LLM outputs with human writing processes.
-Specifically, optimizing prompts with clearer contextual guidance (e.g., "before-text" and intention label definitions) may significantly enhance model performance.
 Finally, the human evaluation process in Section LABEL:sec:appendix:human-eval was determined as exempt from IRB review by the authors’ primary institution, while the data collection using our Chrome extension program was fully approved by the IRB at our institution.
@@
 A.2 Technical Details of System Implementations
-When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel888To prevent privacy concerns, the extension filters out keystroke data from any unauthorized Overleaf projects.
+When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel111To prevent privacy concerns, the extension filters out keystroke data from any unauthorized Overleaf projects.
 Please see Appendix A.2 for more details..
-When each of these actions999Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page.
-occurs, the extension uses ‘diff_match_patch’ package101010https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 6).
+When each of these actions222Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page.
+occurs, the extension uses ‘diff_match_patch’ package333https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 7).
 Then, the extension will send the array along with metadata (e.g., time stamp, author ID, etc.) to the backend server.
@@
 Appendix D ScholaWrite dataset statistics
-Table 6 presents the overall statistics of our ScholaWrite dataset. Table 8 shows the distribution of intention labels per Overleaf project from ScholaWrite.
-Project | 1 | 2 | 3 | 4 | 5 |
-# Authors | 1 (3) | 1 (4) | 1 (3) | 1 (4) | 9 (18) |
-# keystrokes | 14,217 | 5,059 | 6,641 | 8,348 | 27,239 |
-# Words added | 17,387 | 23,835 | 7,779 | 12,448 | 57,511 |
-# Words deleted | 11,739 | 15,158 | 2,308 | 7,621 | 25,853 |
+Table 8 shows the distribution of intention labels per Overleaf project from ScholaWrite.
 | Label | Subsequent label | Probability | |
@@
 | Visual Formatting | Text Production | 0.25 |
-1 | 2 | 3 | 4 | 5 | |
+| 1 | 2 | 3 | 4 | 5 | |
 | Idea Generation | 515 | 130 | 116 | 309 | 3255 |
@@
 | Visual Formatting | 752 | 163 | 43 | 427 | 567 |
-Figures 9 to 12 show several characteristics of human writing process, analyzed from ScholaWrite dataset: (1) Figure 9 - the average Wasserstein distance between each intention distribution and uniform distribution; (2) Figure 10 - distribution of labels over time; (3) Figure 11 for high-level intention distribution over time; and (4) Figure 12 for intention-wise writing activity distribution over time.
-Appendix E More about ScholaWrite Evaluation
-E.1 Training environments
+Figures 10 to 13 show several characteristics of human writing process, analyzed from ScholaWrite dataset: (1) Figure 10 - the average Wasserstein distance between each intention distribution and uniform distribution; (2) Figure 11 - distribution of labels over time; (3) Figure 12 for high-level intention distribution over time; and (4) Figure 13 for intention-wise writing activity distribution over time.
+Appendix E More about Analyses
+Figure 15 shows transition probability between pair of intentions in early stage and later stage.
+Table 9 shows top 6-gram writing intention sequences by session coverage (%)
+| 6-gram Sequence | Percentage |
+| Text Production Clarity Text Production Clarity Text Production Clarity | 3.22 |
+| Clarity Text Production Clarity Text Production Clarity Text Production | 3.22 |
+| Text Production Object Insertion Text Production Object Insertion Text Production Object Insertion | 2.34 |
+| Object Insertion Text Production Object Insertion Text Production Object Insertion Text Production | 2.34 |
+| Idea Generation Idea Organization Idea Generation Idea Organization Idea Generation Idea Organization | 2.05 |
+Appendix F More about Experiments
+F.1 Predicting Next Writing Intention
+F.1.1 Training environments
 BERT & RoBERTa
@@
 Llama
-For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct111111https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library121212https://github.com/unslothai/unsloth.
+For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct444https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library555https://github.com/unslothai/unsloth.
 For the intention prediction task (Sec.
-5.1), here are the hyperparameter setups for the Llama models: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 5; (4) learning rate of ; and (5) AdamW 8-bit optimizer.
+6.1), here are the hyperparameter setups for the Llama models: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 5; (4) learning rate of ; and (5) AdamW 8-bit optimizer.
 Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset.
 For Llama-8B, it took approximately 8 hours on one RTX A5000.
-For the ‘after’ text generation subtask from the iterative self-writing experiment (Sec.
-5.2), we used the following hyperparameter setups for the fine-tuned Llama-8B-Instruct: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 10; (4) learning rate of ; and (5) AdamW 8-bit optimizer.
-Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset.
-Also, it took approximately 12 hours on one NVIDIA L40s.
 GPT-4o
-We used the GPT-4o-2024-08-06 version for the classification and iterative writing experiments. The iterative writing with 100 iterations took approximately 1 hour on each seed.
-E.2 Details About Finetuning Process
-Next Intention Prediction (Sec 5.1)
+We used the GPT-4o-2024-08-06 version.
+F.1.2 Details About Finetuning Process
 The fine-tuning prompt included all possible labels with definitions, task instructions, the “before-text” chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”.
-Differences in prompts were limited to only task instructions (see §E.3 for prompt details).
-To achieve optimal performance while minimizing memory usage, we employed QLoRA Dettmers et al.
-(2024) to fine-tune all linear modules of a 4-bit quantized Llama.
+Differences in prompts were limited to only task instructions (see §G.3 for prompt details).
+To achieve optimal performance while minimizing memory usage, we employed QLoRA (Dettmers et al., 2024) to fine-tune all linear modules of a 4-bit quantized Llama.
 During fine-tuning, we utilized the ‘train_on_response_only’ function provided by the unsloth library, which masks the task instructions, intention label definitions, and “before” text with -100s.
@@
 The model was fine-tuned for one epoch with a batch size of 2, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
-Iterative Self-Writing (Sec 5.2)
-The structure of the “after” text differs slightly to help the model learn where and what edits to make.
-We used the diff_match_patch library to generate a word-level difference array between the “before” and “after” texts.
-Special tokens (<same>, </same>, <del>, </del>, <add>, </add>) were added to the tokenizer, and the difference array was converted into text wrapped with these tokens.
-For example, given a “before” text of ‘‘Bad dog’’ and an “after” text of ‘‘Good dog’’, the difference array would be [(-1, ‘Bad’), (1, ‘Good ’), (0, ‘dog’)].
-This is converted into: <del>Bad </del><add>Good </add><same>dog</same>.
-This transformation was applied only to the “after” text, while the “before” text remained as plain LaTeX text.
-For finetuning, we randomly split the ScholaWrite dataset into training (80%) and testing (20%) sets.
-From each intention label in the test set, we sample up to 300 keystroke entries due to budget constraints.
-For intention prediction, we fine-tune Llama3.1-8B-Instruct on ScholaWrite training set (Llama-8B-SW-pred”) and compare it to baseline models (Llama-8B-Instruct and GPT-4o) from Section 5.1.
-For “after-text” generation, we fine-tune another Llama3.1-8B-Instruct model (“Llama-8B-SW-gen”) using the same dataset, with Llama3.1-8B-Instruct and GPT-4o as baselines.
-The fine-tuning prompt includes task instructions, a verbalizer from human-annotated labels, and “before-text.” While prompts were standardized, task instructions varied by model.
-The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text.
-For fine-tuning, we used QLoRA Dettmers et al.
-(2024) to optimize all linear modules of a 4-bit quantized model while maintaining a small memory footprint.
-Additionally, the embed_tokens and lm_head modules were set as trainable and saved in the final checkpoint.
-To focus training on the response portion (the “after” text), we used the train_on_response_only function, masking the task instructions, verbalizer, and “before” text with -100s.
-This ensures the model learns to generate the "after" text without being influenced by instructional input.
-The model was trained for one epoch with a batch size of 1, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
-During iterative writing, we performed 100 iterations, treating the model’s output under one intention as a single iteration.
-If the intention predicted by the classification model (fine-tuned Llama3.1-8B-Instruct, as described in Sec 5.1) matched the current predicted intention, the model was prompted to edit the text again.
-In this case, the newly generated output was not treated as final output in the iteration, and the iteration did not proceed.
-We moved to the next iteration only when the intention prediction model generated a different intention label than the previous one.
-For model setups, we created three pairs of models for each prediction and generation subtasks as follows:
-LLama-8B-SW: LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Prediction) & LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Generation), independently
-LLama-8B-Instruct: Vanilla LLama-8B-Instruct (Prediction) & Vanilla LLama-8B-Instruct (Generation), independently
-GPT-4o: GPT-4o inference (Prediction) & GPT-4o inference (Generation)
-Also, due to budget constraints, models had different revision strategies.
-Llama-8B-SW-* and Llama-8B-Instruct continued revision until the next predicted intention changed.
-GPT-4o, however, moved to the next iteration regardless.
-We refer to the fine-tuned Llama-8B model as Llama-8B-ScholaWrite (or Llama-8B-SW) and the vanilla model as Llama-8B-Zero.
-E.3 Prompt Templates
-E.3.1 Prediction Prompt for Llama-8B-Zero models
+F.1.3 Prompt Templates
+Prediction Prompt for Llama-8B-Zero models
 “Here are all the possible writing intention labels:
@@
 ## Output:
-E.3.2 Generation Prompt for Llama-8B-Zero models
-“You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
-You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
-Below is the paper you have written so far. Please strictly follow the writing intention given above and insert, delete, or revise at appropriate places in the paper given below.
-Your writing should relate to the paper given below. Do not generate text other than paper content. Do not describe the changes you are making or your reasoning.
-## Input: before_text ”
-E.3.3 Prediction Prompt for Llama-8B-SW models
+Prediction Prompt for Llama-8B-SW models
 “Here are all the possible writing intention labels:
@@
 ## Output:
-E.3.4 Generation Prompt for Llama-8B-Zero models
-“You are a computer science researcher with extensive experience in scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX.
-You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
-Below is the paper you have written so far. Given the paper information below and the corresponding scholarly writing intention, please revise or add to the text to fulfill this writing intention.
-You may insert, delete, or revise text at appropriate places in the given paper.
-Please provide a complete output. Do not generate text that is nonsensical or unrelated to the given paper information.
-## Input: before_text ”
-E.3.5 Classification Prompt for GPT-4o
+Classification Prompt for GPT-4o
 “You are a classifier that identify the most likely next writing intention.
@@
 ## Output:
-E.3.6 Generation Prompt for GPT-4o
+F.1.4 Discussion about Next Intention Prediction Results
+According to Table 6, none of them is reaching 0.7 in F1.
+This is likely due to the intricate nature of the task.
+The model is asked to predict the next intention by only giving the before text.
+In the annotation task, the annotator labels the data by looking through multiple consecutive before and after text pairs to determine the current intention rather than looking at the current text to predict the next intention.
+Another reason is that the next intention chosen by the author does not necessarily mean that it is the only correct intention.
+We also noticed that BERT and RoBERTa perform much better than the vanilla Llama-8b-Instruct.
+This is likely due to Llama 8b being under-trained as they have a larger size of parameters, while we fine-tuned it on our data for only one epoch.
+F.2 Output Alignment across Writing Sessions
+F.2.1 Prompt Templates
+Single-intention Session Output Generation
+You are an expert writing assistant.
+Your role is to improve user-provided text according to specific criteria.
+The input will be LaTeX text.
+Always output only the improved LaTeX text, with no explanations or notes.
+Improve the following LaTeX text according to this criteria:
+- {intention}: {intention definition}
+Text to improve:
+{before_text}
+Multi-intention Session Output Generation
+You are an expert writing assistant.
+Your role is to improve user-provided text according to specific criteria.
+The input will be LaTeX text.
+Always output only the improved LaTeX text, with no explanations or notes.
+Improve the following LaTeX text according to these criteria:
+{list of intentions and their definitions}
+Text to improve:
+{before_text}
+F.2.2 Results
+Figure 16 shows the relationship between a single-intention session duration and alignment score (either Levenshtein ratio or BERTScore-F1) for two models, GPT-5 and Qwen-2.5-14B.
+Figure 17 shows the relationship between a single-intention session duration and alignment score (either Levenshtein ratio or BERTScore-F1) for two models, GPT-5 and Qwen-2.5-14B across different intention categories.
+Figure 18 shows the relationship between number of writing intentions in a multi-intention session and alignment score (either Levenshtein ratio or BERTScore-F1) for two models, GPT-5 and Qwen-2.5-14B.
+Appendix G More about ScholaWrite Evaluation
+We envision ScholaWrite as a valuable resource for training language models and improving future writing assistants for scholarly writing.
+To evaluate its usability, we conducted experiments training LLMs to mimic the complex, non-linear writing processes of human scholars.
+Iteratively generating scholarly writing actions from scratch (called Iterative Self-Writing), mirroring the human writing process: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
+Setups
+During iterative self-writing (Figure 19), a fine-tuned model processes LaTeX-formatted seed document (as “before-text”) with a context prompt to predict the next intention, then revises the text (“after-text”) accordingly given prompt.
+The revised document then serves as the new seed for the next iteration.
+This process repeats until a set iteration limit (e.g., 100) is reached.
+All models use the same train (80%)-test (20%) split across experiments.
+Figure 23 for training details.
+We fine-tune Llama3.1-8B-Instruct (Llama-8B-SW) and compare it to vanilla Llama-8B-Instruct (Llama-8B-Zero) and GPT-4o.
+Also, seed documents were derived from LaTeX-formatted abstracts of four award-winning NLP papers on diverse topics (Zeng et al., 2024; Lu et al., 2024; Du et al., 2022a; Etxaniz et al., 2024), as shown in Appendix Listings LABEL:table:seed-entry-johnny-LABEL:table:seed-entry-latxa.
+Metrics
+We evaluated lexical diversity (unique-to-total token ratio), topic consistency (cosine similarity between seed and final output), and intention coverage (unique writing intentions used over 100 iterations).
+For human evaluation, three native English speakers with LaTeX expertise assessed outputs from Llama-8B-Zero and Llama-8B-SW on accuracy (alignment with predicted intention), alignment (similarity to human writing), fluency(grammatical correctness), coherence(logical structure), and relevance (connection to the seed document) - Refer to Appendix LABEL:sec:appendix:human-eval.
+Accuracy was judged per iteration, while other metrics used pairwise comparisons.
+Inter-annotator agreement (IAA)666The IAA scores are 0.84 (SW) and 0.76 (Zero) for accuracy, all 100% for alignment, fluency, coherence, and 49.8% (SW) and 100% (Zero) for relevance.
+was measured using Krippendorff’s alpha for accuracy and percentage agreement for others.
+Llama-8B-SW
+Llama-8B-Zero
+GPT-4o
+| Iter. | Llama-8B-SW | Llama-8B-Zero | GPT-4o |
+[…]
+[…]
+[…]
+Results
+Figure 20 shows that Llama-8B-SW consistently produced the most lexically diverse words, generated the most semantically aligned topics (Seeds 1 & 2), and covered the most writing intentions (except Seed 3).
+These results underscore the value of ScholaWrite in improving scholarly writing quality generated by language models.
+However, our human evaluation (Figure 21) revealed that Llama-8B-SW generated less human-like writing, in terms of fluency and logical claims.
+It also struggled with generating texts aligned with the predicted intentions.
+See Appendix Tables LABEL:table:human-eval-seed1 to LABEL:table:human-eval-seed4 for more details.
+Despite the weaknesses, Llama-8B-SW still produced more relevant content (Seed 2), which aligns with topic consistency trends in Figure 20, highlighting the usefulness of ScholaWrite dataset in certain contexts.
+Moreover, Llama-8B-SW exhibited the most human-like writing activity patterns over time (Figure 22), which frequently switches between implementation and revision and covers all three high-level processes.
+Llama-8B-Zero and GPT-4o tend to remain in a single high-level stage throughout all 100 iterations of self-writing (see Appendix LABEL:fig:writing-step-broad-all-model and LABEL:fig:writing-step-intention-all-model for details).
+Compared to Appendix Figure 13, which depicts frequent transitions across all three stages in an early draft (e.g., the first 100 steps), Llama-8B-SW most closely replicates human writing behaviors in iterative writing tasks.
+These findings reinforce the potential of ScholaWrite in helping LLMs emulate human scholarly writing processes.
+G.1 Training environments
+BERT & RoBERTa
+We fine-tuned BERT and RoBERTa with the following hyperparameter setups: (1) a learning rate of ; (2) training batch size per device of 8; (3) evaluation batch size per device of 8; (4) the number of training epochs of 10; and (5) a weight decay of .
+For each model, it took approximately 3.5 hours on one NVIDIA RTX A6000.
+Llama
+For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct777https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library888https://github.com/unslothai/unsloth.
+For the ‘after’ text generation subtask from the iterative self-writing experiment, we used the following hyperparameter setups for the fine-tuned Llama-8B-Instruct: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 10; (4) learning rate of ; and (5) AdamW 8-bit optimizer.
+Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset.
+Also, it took approximately 12 hours on one NVIDIA L40s.
+GPT-4o
+We used the GPT-4o-2024-08-06 version. The iterative writing with 100 iterations took approximately 1 hour on each seed.
+G.2 Details About Finetuning Process
+The structure of the “after” text differs slightly to help the model learn where and what edits to make.
+We used the diff_match_patch library to generate a word-level difference array between the “before” and “after” texts.
+Special tokens (<same>, </same>, <del>, </del>, <add>, </add>) were added to the tokenizer, and the difference array was converted into text wrapped with these tokens.
+For example, given a “before” text of ‘‘Bad dog’’ and an “after” text of ‘‘Good dog’’, the difference array would be [(-1, ‘Bad’), (1, ‘Good ’), (0, ‘dog’)].
+This is converted into: <del>Bad </del><add>Good </add><same>dog</same>.
+This transformation was applied only to the “after” text, while the “before” text remained as plain LaTeX text.
+For finetuning, we randomly split the ScholaWrite dataset into training (80%) and testing (20%) sets.
+From each intention label in the test set, we sample up to 300 keystroke entries due to budget constraints.
+For intention prediction, we fine-tune Llama3.1-8B-Instruct on ScholaWrite training set (Llama-8B-SW-pred”) and compare it to baseline models (Llama-8B-Instruct and GPT-4o) from Section 6.1.
+For “after-text” generation, we fine-tune another Llama3.1-8B-Instruct model (“Llama-8B-SW-gen”) using the same dataset, with Llama3.1-8B-Instruct and GPT-4o as baselines.
+The fine-tuning prompt includes task instructions, a verbalizer from human-annotated labels, and “before-text.” While prompts were standardized, task instructions varied by model.
+The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text.
+For fine-tuning, we used QLoRA (Dettmers et al., 2024) to optimize all linear modules of a 4-bit quantized model while maintaining a small memory footprint.
+Additionally, the embed_tokens and lm_head modules were set as trainable and saved in the final checkpoint.
+To focus training on the response portion (the “after” text), we used the train_on_response_only function, masking the task instructions, verbalizer, and “before” text with -100s.
+This ensures the model learns to generate the "after" text without being influenced by instructional input.
+The model was trained for one epoch with a batch size of 1, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
+During iterative writing, we performed 100 iterations, treating the model’s output under one intention as a single iteration.
+If the intention predicted by the classification model (fine-tuned Llama3.1-8B-Instruct, as described in Sec 6.1) matched the current predicted intention, the model was prompted to edit the text again.
+In this case, the newly generated output was not treated as final output in the iteration, and the iteration did not proceed.
+We moved to the next iteration only when the intention prediction model generated a different intention label than the previous one.
+For model setups, we created three pairs of models for each prediction and generation subtasks as follows:
+LLama-8B-SW: LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Prediction) & LLama-8B-Instruct fine-tuned on ScholaWrite dataset (Generation), independently
+LLama-8B-Instruct: Vanilla LLama-8B-Instruct (Prediction) & Vanilla LLama-8B-Instruct (Generation), independently
+GPT-4o: GPT-4o inference (Prediction) & GPT-4o inference (Generation)
+Also, due to budget constraints, models had different revision strategies.
+Llama-8B-SW-* and Llama-8B-Instruct continued revision until the next predicted intention changed.
+GPT-4o, however, moved to the next iteration regardless.
+We refer to the fine-tuned Llama-8B model as Llama-8B-ScholaWrite (or Llama-8B-SW) and the vanilla model as Llama-8B-Zero.
+G.3 Prompt Templates
+G.3.1 Generation Prompt for Llama-8B-Zero models
+“You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
+You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
+Below is the paper you have written so far. Please strictly follow the writing intention given above and insert, delete, or revise at appropriate places in the paper given below.
+Your writing should relate to the paper given below. Do not generate text other than paper content. Do not describe the changes you are making or your reasoning.
+## Input: before_text ”
+G.3.2 Generation Prompt for Llama-8B-Zero models
+“You are a computer science researcher with extensive experience in scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX.
+You currently want to ‘‘put the verbalizer of the predicted intention label’’ (e.g., ‘‘initially create sections and sub-level structures’’ if the predicted label was section planning).
+Below is the paper you have written so far. Given the paper information below and the corresponding scholarly writing intention, please revise or add to the text to fulfill this writing intention.
+You may insert, delete, or revise text at appropriate places in the given paper.
+Please provide a complete output. Do not generate text that is nonsensical or unrelated to the given paper information.
+## Input: before_text ”
+G.3.3 Generation Prompt for GPT-4o
 “ You are a computer science researcher with extensive experience of scholarly writing. Here, you are writing a research paper in natural language processing using LaTeX languages.
@@
 ## Output:
-E.4 Seed Documents for Iterative Self-Writing
-We present the remaining three seed documents that we used for the iterative self-writing experiments, as shown in Listings LABEL:table:seed-entry-johnny to LABEL:table:seed-entry-latxa.
+G.4 Seed Documents for Iterative Self-Writing
+We present the four seed documents that we used for the iterative self-writing experiments, as shown in Listings LABEL:table:seed-entry-johnny to LABEL:table:seed-entry-latxa.